A Complete Bibliography of Publications in Cryptography and Communications: Discrete Structures, Boolean

Functions and Sequences

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

29 July 2022 Version 1.12

Title word cross-reference

(−1) [519]. (1 − 2u^{3}) [185]. (4n, 2, 4n, 2n) [225]. (k, R, 1) [428]. (n, n − 1) [240]. (n, n − 2) [240]. (p^{n}, p^{n}, p^{n}, 1) [35]. (r, ≤ 2) [15]. (σ, δ) [453]. (x^{pm} − x + δ)^{s} + x^{pm} + x [237]. 1 [273, 174]. 105 [129]. 128 [288]. 16 [381]. 1\frac{1}{2} [357]. 2 [244, 243, 142, 203, 273, 182, 231, 293, 403, 438]. 22 [232]. 24k + 10 [310]. 2^{k} [172, 430]. 2^{n} [138, 191]. 2p^{n} [521]. 2p^{s} [439]. 3 [236]. 30 [68]. 4 [240, 432, 181, 179, 163]. 4k + 2 [310]. 4N [488]. 4 × 4 [89]. 5 [173]. 6 [277, 26, 186]. 8 [240]. 8q [509]. ^{2} [468]. a [209]. c [452, 490]. C^{∗} [301]. C_{n} × Q_{8} [225]. cx + Tr_{q^{l}/q}(x^{a}) [235]. d [404]. ℓ [143]. ℓ^{t}p^{s} [104]. f^{d} [373]. f^{v} [69]. F_{p} [292]. F_{p}[u, v]/⟨u^{2} − 1, v^{3} − v, uv − vu⟩ [292]. G [474]. G(X)^{k} − L(X) [451]. GF(2^{k}) [18]. GF(2^{n}) [44]. GF(l) [124]. j [467]. k [148, 7]. l [10]. L_{1} [272]. L_{1}(x^{3}) + L_{2}(x^{9}) [272]. L_{2} [272]. m [188, 385]. M_{2}(F_{2} + uF_{2}) [265]. F_{2^{8}} [470]. F_{2^{n}} [412]. F_{2} × (F_{2} + vF_{2}) [457, 444]. F_{5^{n}} [291]. F_{p} [398, 502]. F_{p} + uF_{p} [195]. F_{p}[u]/⟨u^{m} − u⟩ [224]. F_{p}[u]/⟨u^{4} − u⟩ [185]. F_{p^{m}} [242]. F_{p^{m}} + uF_{p^{m}} + u^{2}F_{p^{m}} [439]. F_{q} + uF_{q} [360]. F_{q^{2}} + uF_{q^{2}} [513]. Fp^{2m} [237]. P^{3} [203]. Q [523]. R [523]. Z_{2^{m}} [127]. Z_{2}Z_{2}[u, v] [388]. Z_{4} [184, 160, 170, 174]. Z_{4} + uZ_{4} [172, 430, 453]. Z_{4}[v]/⟨v^{2} + 2v⟩ [377]. Z_{p^{2}} + uZ_{p^{2}} [529]. Zpe [282, 330]. Z_{p^{k}} [350]. Z_{p}Z_{p^{s}} [374]. Z/(p^{e}q) [142]. B_{j,k} [474]. max(R) = 3 [428]. N [386, 318]. N − 2 [318]. n = \frac{q^{2m}−1}{2} [369]. \overline{2} [181]. \overline{3} [236]. p [502, 276, 260, 538, 427, 343, 150, 165]. p^{n} [521, 148]. pq [184, 38]. pq^{2} [456]. q [244, 262, 326, 128]. r [333]. S [421, 462]. t [424, 492, 102, 241, 447, 533]. x + x^{2^{l}} + ⋯ + x^{2^{ml}} = a [412]. x^{2^{l}+1} + x + a [18]. x^{r}h(x^{s}) [289]. Z [482, 540]. Z_{4}[u]/⟨u^{k}⟩ [193]. Z_{r} = r [316].

-additive [388, 374]. -adic [231, 293, 403, 438]. -ary [244, 326, 276, 260, 538, 427, 343, 150, 165]. -boxes [421]. -byte [232]. -codes [474]. -complementary [540, 482]. -complementary/complementary [540]. -constacyclic [185]. -designs [424, 492, 533]. -differential [452, 490, 519]. -divisibility [243]. -error [148, 7]. -Feistel [209]. -fold [182, 102]. -functions [240]. -generator [174]. -identifying [15]. -invariant [467]. -Linearized [502]. -packings [428]. -PD [462]. -periodic [456]. -polynomial [128]. -QAM [381]. -relative [225, 35]. -separable [181]. -sequences [143, 188, 385, 170, 10]. -Skew [453]. -strongly [236]. -th [333, 386]. -to- [273]. -transform [262]. -tuple [241]. -uniform [240, 432, 179]. -variable [277]. -wise [447].

16 [294].

2 [329]. 2-Adic [488, 509]. 2-designs [461, 511]. 256 [216].

64 [91].

7 [75].

8 [294]. 8-bit [475]. 80 [86].

9 [430].

abelian [455, 324, 45]. absolute [304].

Absorbing [342]. accelerator [63]. access [194]. ACD [532]. achieving [160]. active [152]. addition [138, 491, 191, 421]. additive [423, 388, 374]. Adic [488, 509, 231, 293, 403, 438]. adversary [39]. AEAD [531]. AES [99, 250, 4, 471, 441]. AES-like [250, 99, 4, 441]. Affine [296, 494, 286, 18, 335, 511, 524]. affine-invariant [511]. again [415]. against [86, 257, 114]. algebra [111]. Algebraic [440, 240, 357, 92, 138, 70, 481, 46, 88, 347, 61, 230]. algorithm [144, 273, 232, 531, 36]. algorithms [468, 137, 512]. Almost [408, 322, 427, 459, 335, 111, 534, 482, 283, 231, 347]. alphabet [244]. alphabets [233]. Analysis [254, 141, 300, 475, 458, 252, 51, 59, 209, 8, 189, 211]. analytic [33]. aperiodic [518, 314]. APN [546, 37, 394, 517, 414, 345, 494, 321, 273, 346, 495, 135, 416, 272, 508, 131, 519]. Application [144, 507, 42, 250, 460, 9, 490]. applications [526, 278, 132, 344, 271, 413, 489, 475, 332, 105, 76, 378, 21, 478, 273, 270, 245, 547, 305, 183, 59, 146, 154, 373, 513, 91]. Applying [56, 500]. approach [41, 112, 125, 33]. approximate [454, 528]. approximation [191]. arbitrary [269]. arising [354]. arithmetic [399]. array [123]. arrays [80, 202, 111, 113, 65]. ary [244, 326, 276, 260, 538, 427, 343, 150, 165]. Ashikhmin [476]. assisted [499, 498, 206]. associated [243, 448]. asymmetric [477, 372, 371]. Asymptotic [19, 82]. Asymptotically [181, 323, 113, 410, 436, 374]. Asynchronous [290]. attack [98, 216, 458, 12, 96, 109, 91, 473]. Attacking [527]. Attacks [106, 54, 90, 53, 255, 152, 99, 56, 101, 86, 75, 257, 5, 47, 138, 55, 209, 251, 171, 8]. authenticated [147]. Authentication [227, 147, 23, 102, 182, 87]. autocorrelation [408, 233, 202, 117, 509, 387, 283, 231, 403, 488, 38, 382]. Autocorrelations [10]. automata [41, 274]. automatic [258, 539]. automorphisms [417]. Average [134].

Backtracking [206]. Backtracking-assisted [206]. balanced [117, 493, 283, 416, 347, 239]. Barg [476]. base [484]. Based [227, 469, 184, 152, 496, 101, 180, 84, 105, 188, 106, 274, 162, 473, 189, 472, 36, 230, 319, 487]. bases [25, 26, 454, 27, 77]. basic [80]. BCH [369, 367, 530, 375, 533]. beautiful [31]. behavior [517, 452]. behaviour [134]. Bent [158, 116, 284, 353, 418, 160, 64, 491, 504, 417, 465, 352, 276, 260, 350, 538, 136, 162, 389, 354, 45, 220, 196, 343, 362, 201, 131, 165, 522, 379, 433]. best [191]. Beth [152]. between [363, 508]. beyond [215, 257, 249]. beyond-birthday-bound-secure [215]. beyond-the-birthday [249]. beyond-the-birthday-bound [257]. BGLS [208]. bias [146]. biased [175]. biases [11]. bifix [78, 93]. big [273]. bijective [137, 89, 421]. Binary [315, 222, 393, 318, 376, 462, 521, 214, 247, 332, 411, 64, 148, 387, 525, 401, 535, 361, 476, 302, 231, 403, 383, 329, 523, 488, 456]. biquadratic [301]. birthday [215, 257, 249]. bit [475, 333, 212]. bits [345, 97]. BKW [141]. black [472]. Block [71, 147, 99, 249, 528, 503, 441]. blockchain [217]. blockcipher [106]. blockcipher-based [106]. BLS [208]. boolean [464, 140, 243, 107, 347, 43, 50, 356, 132, 344, 271, 413, 394, 489, 277, 420, 19, 261, 144, 126, 82, 151, 175, 92, 47, 460, 70, 351, 275, 402, 493, 320, 46, 524, 156, 88, 61, 325]. boomerang [434, 520, 537]. bordered [366, 409]. both [100]. bound [215, 257, 160, 249, 518, 320, 476, 293, 362]. boundary [112]. Bounds [72, 267, 236, 328, 356, 247, 290, 23, 312, 6, 326, 407, 15, 524, 288]. box [472]. boxes [48, 95, 475, 137, 542, 313, 178, 441, 274, 298, 89, 421]. BP [123]. BP-XOR [123]. Bruijn [219, 241]. burn [254]. burn-in [254]. burst [372, 371]. butterfly [345]. byte [232, 339, 372, 371]. bytes [11].

CAR30 [68]. Carlitz [478]. Cartesian [391]. cases [112, 14]. CAST [216]. CAST-256 [216]. Categorizing [470]. CCZ [363, 459, 495]. CCZ- [363]. CCZ-equivalence [495]. CCZ-inequivalence [459]. CDMA [384]. Cellular [274]. central [35]. Certain [399, 547]. chain [545, 501, 317, 234]. Changing [346]. channel [255, 458, 290, 472]. Character [197]. characterising [171]. Characteristic [450, 246, 418, 269, 415, 203, 348, 235, 279, 534, 69, 308, 131]. Characterization [107, 398]. characterizations [23]. Characters [490, 351]. Cheating [65]. Cheating-immune [65]. Chebyshev [544]. check [144]. chi [256]. chi-square [256]. Chinese [125]. Chudnovsky [468]. cipher [147, 68, 173, 215, 75, 119, 12, 13, 83, 60, 88, 49, 58]. ciphers [99, 250, 4, 56, 63, 252, 257, 5, 249, 55, 9, 441, 96, 294, 87, 189, 487]. circuits [278]. Circulant [22, 483, 214, 105, 312, 360]. circulant-like [105]. class [259, 180, 348, 134, 186, 504, 479, 103, 397, 511, 535, 350, 331, 231, 242, 403, 375, 221, 196, 362, 308, 446, 164, 166, 311, 487]. Classes [515, 464, 66, 394, 414, 353, 184, 126, 285, 35, 177, 199, 182, 280, 401, 264, 266, 352, 276, 7, 334, 486, 476, 190, 431, 291, 165, 229, 438]. classical [158, 303]. Classification [220, 541, 377]. closed [334]. co [304]. co-trace [304]. coalitions [114]. cocyclic [29]. code [122, 244, 469, 450, 336, 361, 405, 383, 288]. code-based [469]. codebooks [160, 323, 436]. coded [40]. Codes [437, 507, 306, 457, 444, 364, 498, 376, 541, 292, 327, 78, 172, 328, 463, 342, 400, 445, 93, 159, 193, 377, 430, 67, 85, 462, 181, 470, 324, 157, 223, 492, 496, 545, 366, 474, 180, 536, 247, 84, 125, 332, 357, 286, 409, 224, 341, 369, 315, 310, 461, 177, 130, 411, 23, 501, 428, 335, 312, 423, 186, 547, 479, 529, 326, 503, 74, 102, 103, 135, 161, 367, 396, 422, 530, 182, 178, 238, 280, 401, 511, 535, 222, 265, 264, 266, 128, 477, 442, 453, 162, 334, 309, 510, 410, 407, 476, 123, 15, 317, 448, 331, 339, 372, 371, 185, 447]. codes [14, 440, 104, 155, 127, 228, 195, 268, 388, 359, 360, 480, 532, 234, 439, 242, 375, 221, 190, 176, 163, 319, 431, 446, 497, 174, 543, 149, 164, 443, 533, 150, 165, 513, 522, 287, 166, 311, 374, 167, 236, 449, 65, 499]. codewords [74]. coding [526, 255, 270, 245, 281]. coefficients [92]. coherence [395]. Cohn [197]. collection [329]. colliding [232]. collision [98, 232]. combination [53]. Combinatorial [424, 23, 73, 8]. Comments [61]. Commun [430]. communication [173, 281]. Communications [1]. commutative [34, 547, 234]. compartmented [194]. Complementary [322, 364, 376, 81, 341, 315, 482, 425, 228, 540, 77, 381]. complementary-dual [228]. complementary/complementary [540]. Complete [377, 99, 177, 264, 190, 311, 167, 290, 479, 166]. completed [504]. complex [25, 24]. Complexity [267, 505, 468, 356, 277, 420, 184, 117, 198, 521, 509, 2, 118, 484, 542, 6, 148, 337, 7, 187, 69, 231, 293, 403, 386, 523, 124, 318, 488, 438, 89, 406]. component [435]. component-wise [435]. components [218, 418]. composition [370, 329]. Compositional [289]. compressed [395, 319, 282]. compressing [110, 330]. computational [469]. Computing [304, 92, 69, 438, 146]. concatenated [341]. concatenation [313]. concurrent [152]. Conditional [129]. Conditions [116]. conflict [445]. congruential [527]. conjecture [546, 112, 358]. conjectures [133, 387]. connection [10]. consisting [2]. constacyclic [185, 104, 155, 359, 439, 375]. constant [503, 510]. constraints [113]. construct [239, 200]. constructed [259, 153, 313, 543]. Constructing [131, 465]. Construction [219, 462, 116, 501, 322, 96, 127, 220, 38, 78, 463, 48, 324, 223, 341, 380, 528, 102, 161, 390, 183, 162, 510, 77, 149]. Constructions [240, 298, 168, 499, 498, 395, 72, 366, 536, 290, 409, 503, 422, 70, 518, 323, 477, 136, 493, 368, 73, 319, 436, 488, 287, 236, 449, 200, 381]. containing [530]. context [90]. continued [523]. conversely [284]. convolutional [309]. cookie [251]. Correcting [430, 180, 306, 326, 477, 361, 331, 339, 372, 371]. Correction [457, 499, 270, 372]. Correlation [53, 262, 54, 98, 81, 19, 506, 385, 118, 333, 153, 518, 113, 170, 398, 379, 314]. correlation-immune [19, 398]. correlations [393]. cost [458]. Counting [126, 171, 134, 33]. covering [328, 145, 288]. CRC [227]. criteria [4, 139]. criterion [29]. cross [78, 93, 393]. cross-bifix-free [78, 93]. Cryptanalysis [207, 531, 60, 294, 267, 487, 129, 296, 46, 253, 61]. Cryptogr [430]. Cryptographic [42, 175, 88, 421, 256, 278, 137, 64, 422, 183, 146, 96]. Cryptographically [227, 105, 139]. Cryptography [295, 101, 204, 1]. cryptosystems [301]. cube [56, 320]. CubeHash [59]. cubic [50, 66, 402]. cumulative [65]. curves [451, 505, 98, 527, 467]. Cusick [112]. cut [109]. cycle [269, 373]. cycles [58]. Cyclic [193, 186, 265, 234, 292, 159, 377, 324, 223, 545, 125, 357, 177, 501, 529, 103, 535, 477, 453, 228, 359, 242, 163, 497, 174, 150, 374, 449]. cyclotomic [184, 521, 186, 337, 486, 154, 438]. cyclotomy [198, 480, 163].

D [244]. Data [267, 257]. deception [102]. Deciding [514]. Decimated [385]. decision [302]. decodable [39]. decoding [423]. Decomposition [297, 340]. decompositions [486]. decorrelation [90]. defined [34]. degree [240, 404]. degree- [404]. degrees [123]. delay [183]. density [28]. depth [278]. derivatives [433]. derived [43, 505, 352, 539, 221, 282, 456]. descendants [307]. Design [198, 252, 441, 20, 46, 61]. Designed [367]. Designing [63]. designs [424, 492, 461, 528, 511, 533]. detection [440]. determinants [9]. determine [7]. Determining [148]. Deterministic [319, 395]. developments [128]. deviation [209]. devices [466]. diagrams [302]. Difference [322, 225, 408, 307, 35, 455, 239]. differences [355]. Differential [542, 520, 267, 370, 429, 129, 86, 415, 321, 55, 534, 96, 537, 171, 189, 452, 490, 139, 519]. differentially [240, 179]. differing [67, 85]. diffusion [183]. digit [246]. digit-sum [246]. digital [51]. digits [484]. dimension [126, 26, 454, 503, 510]. dimensional [410, 448]. dimensions [414, 27]. Ding [337, 124, 438]. Ding-Helleseth [124]. direct [390]. discrete [333]. distance [336, 411, 396, 268, 242, 497, 150]. distance-optimal [396]. distances [529, 367]. distinctness [426, 142]. distinguisher [250]. distinguishing [75, 5, 47, 12]. distribution [450, 442, 241, 231, 242, 221, 170, 282]. distributions [177, 97]. divisibility [112, 243]. DNA [445]. domain [119]. Double [409, 106, 372, 371, 360]. doubly [113]. DryGASCON [531]. dual [444, 364, 376, 496, 366, 474, 536, 409, 315, 310, 547, 417, 390, 530, 352, 317, 104, 127, 228, 360, 220, 287, 457]. dual-containing [530]. duality [485]. duals [136, 446, 497].

EA [363, 414, 495, 514, 44, 452]. EA- [495].

EA-classes [414]. EA-equivalence [514, 452]. EA-equivalences [363]. EA-equivalent [44]. EAQECCs [507, 513]. Eastman [197]. edge [85]. Editorial [132, 344, 271, 295, 413, 489, 1, 94, 204, 378, 20, 76, 52, 248]. Efficient [210, 486, 46, 61]. eigenanalysis [151]. eight [337, 405]. elements [158, 304, 532, 282]. elliptic [98, 527, 467]. embedded [466]. embedding [516]. encoding [123]. encryption [207, 147, 213, 87]. engineering [99]. Enhanced [139]. ensemble [130]. entanglement [499, 498, 513]. entanglement-assisted [499, 498]. entropy [217]. Enumeration [228, 19]. enumerators [479, 264, 190, 166, 311, 167]. equal [102]. Equivalence [66, 35, 126, 321, 495, 514, 452, 24, 508]. equivalences [363]. equivalent [494, 44]. Error [300, 39, 180, 6, 326, 148, 477, 7, 361, 331]. error-correcting [326, 477]. errors [145, 306, 339, 372, 371]. Espresso [173]. estimate [521, 524]. Euclidean [437]. Euler [456]. evaluation [471, 472]. even [310, 79, 235, 279, 406]. Exact [243, 231]. existence [120]. Exotic [24]. expander [210]. Expansion [187, 386]. Explicit [267, 330]. exponential [243, 397]. exponents [259, 285, 199, 279, 305, 222, 338, 291]. Extended [114, 98, 335]. extender [119]. external [455].

F [523]. Factorization [9, 547, 302, 419]. families [307, 459, 384, 385, 455, 397, 136, 113, 508, 443, 433]. family [202, 506, 369, 396, 83, 96, 8, 347, 329, 497]. fast [54, 481, 46, 171, 61]. Faster [232]. Fault [51, 473, 255, 101, 441, 96, 189]. fault-resilient [441]. FCSR [41, 6]. FCSRs [183]. feedback [230]. feedbacks [57, 121]. Feedforward [365]. Feistel [42, 209]. Fermat [43, 203]. few [37, 280, 162, 393, 164, 443]. Fibonacci [180]. field [304, 297]. fields [451, 327, 468, 269, 223, 198, 544, 125, 478, 9, 479, 235, 289, 279, 367, 442, 187, 368, 407, 155, 228, 308, 131, 168]. filtering [241]. Finding [62]. finite [451, 468, 342, 304, 269, 223, 198, 544, 125, 478, 501, 9, 547, 479, 235, 289, 279, 367, 442, 187, 297, 368, 407, 155, 228, 234, 308, 131, 168]. FIRE [99]. Five [266, 343, 497]. flat [325]. flexible [233, 396]. flower [400]. fold [102, 182]. fool [472]. form [451, 92, 235, 289, 237]. Formal [485]. formally [390]. forms [167]. formula [172, 430, 478]. formulae [228]. four [447, 532, 150, 165]. four-weight [447]. Fourier [398, 33]. Fourier-analytic [33]. fractions [523]. frames [392]. framework [189]. free [78, 445, 93, 130, 110]. Frequency [314, 391, 384, 188, 518, 40, 365, 229, 200, 515, 516]. Frequency-hopping [314, 391, 188, 229, 200, 515, 516]. Frobenius [353, 409]. full [404]. function [464, 144, 435, 92, 484, 44, 534, 179, 88, 452, 490]. functional [544]. Functions [116, 158, 240, 43, 218, 429, 50, 140, 37, 356, 66, 132, 344, 271, 413, 394, 489, 414, 277, 420, 19, 261, 243, 353, 418, 494, 126, 82, 157, 424, 151, 415, 175, 17, 459, 321, 262, 160, 47, 340, 478, 273, 64, 460, 346, 495, 134, 491, 504, 417, 471, 107, 103, 135, 422, 465, 534, 70, 352, 276, 481, 260, 350, 538, 136, 162, 351, 275, 389, 402, 493, 416, 16, 354, 46, 284, 45, 171, 524, 156, 88, 440, 370, 220, 139, 196, 347, 343, 362, 272, 508, 61, 398, 201, 325, 436, 543, 131, 165, 522, 519, 239]. functions [379, 433, 168]. Further [136, 108, 121, 338, 449].

G [173]. Gabidulin [423]. Galois [507, 487, 512]. gaps [73]. Gaussian [448]. General [116, 491, 39, 253]. generalisations [411]. generalised [345, 379]. generalizations [301, 75]. Generalized [233, 116, 350, 154, 200, 184, 198, 521, 536, 125, 461, 160, 423, 186, 74, 337, 401, 538, 351, 389, 537, 14, 189, 268, 480, 170, 174, 438, 512, 42]. generate [336]. generated [365]. generating [214, 340, 273]. generation [81, 137, 486]. generation/correlation [81]. generator [527, 174]. generators [53, 241, 365]. Generic [209, 162, 149, 239]. genetic [137]. genus [505]. geometries [342]. GGHN [13]. girth [299, 405]. giving [353]. Golay [496]. Gold [494, 490]. Good [245, 172, 430, 137, 374, 314, 270]. Goppa [159]. grain [252, 96, 129]. grain-like [252]. graph [130]. graphical [84]. graphs [67, 85, 299, 544, 210, 312, 354]. gray [93, 224]. greedy [169]. grid [15]. group [366, 409, 417]. groups [455, 120]. Grover [144]. GRS [507]. Guest [94, 20, 52, 248].

Hadamard

[29, 25, 307, 462, 31, 21, 425, 30, 32, 24, 28, 33]. half [143]. half- [143]. Hamming [62, 385, 401, 518, 314]. hardware [183, 473]. hash [531, 106, 16]. Hashing [467]. having [390]. held [492]. Helleseth [133, 337, 124, 438]. Hermitian [541, 530, 513]. High [55, 240, 198]. highly [48, 440]. hit [188, 200, 391, 515]. Hitag2 [88]. homomorphic [207]. hop [525]. hopping [290, 384, 188, 518, 229, 200, 314, 391, 515, 516]. Horizontal [98]. hull [448]. Hulls [437, 507]. hyper [389, 196]. hyper-bent [389, 196]. hyperelliptic [505]. hyperplane [203]. hypotheses [253].

IDEA [216]. ideal [119, 241, 113].

idempotents [359]. identification [152]. identifiers [336]. identifying [67, 85, 130, 15]. identity [152]. identity-based [152]. If [345]. II [203]. illustration [471]. Image [355]. images [224]. immune [19, 262, 107, 398, 65]. immunity [70, 481, 88, 347]. implementation [141, 46, 61]. implementations [95, 471]. Improved [119, 512, 96]. Improving [524]. index [357, 340, 317]. indicator [491]. individual [97]. inducing [504]. inductive [463]. inequivalence [29, 459]. infinite [353, 136, 15, 433]. infinitely [394]. Influence [138]. information [114, 211]. initialisation [87]. initialization [58]. injection [255]. injective [330]. Injectivity [110, 282]. input [275]. inputs [175]. inserting [503]. instances [300]. Integer [339, 372, 371, 10]. integers [110, 270, 245]. integral [216, 250]. integrated [63]. intercept [153]. interleaved [118]. intersection [447]. invariant [467, 511]. Invariants [495, 514]. Inverse [435, 44, 534, 179, 452]. inverses [289]. invertible [214]. Investigation [475]. involutions [465, 352, 368]. IPM [470]. Irreducibility [227, 203]. irreducible [219, 359]. Issue [271, 295, 1, 466, 526, 132, 344, 413, 489, 20, 76, 204, 378]. iterated [90]. iterations [16]. iterative [340].

Jacobi [293]. Joint [2, 148]. Jump [525].

Karatsuba [205]. Kasami [17]. KATAN32 [91]. KATAN32/48/64 [91]. key [86, 301, 232, 12, 11, 253, 399]. keystream [11]. Kim [494]. Kim-type [494]. kind [176]. king [15]. KISS [212]. Klein [286]. Kloosterman [426, 221, 196, 343, 303]. known [414, 508, 28].

Large [384, 153, 299, 268]. largest [320]. Latin [27]. lattice [527, 36, 230]. lattice-based [36, 230]. LCD [541, 247, 501, 547, 367, 530, 407, 317, 359, 360, 532, 513]. LDPC [405]. leakage [100]. leakage-resilient [100]. learning [472]. least [333]. Lee [529]. Legendre [153]. Lempel [197]. length [172, 193, 430, 181, 509, 369, 310, 306, 106, 361, 104, 540, 439, 57, 36, 121, 170, 163, 288, 38, 236, 406, 381]. length/space/time [410]. lengths [79]. Levenshtein [160]. LFSR [269, 6]. LFSRs [183]. Li [112]. lifted [122]. lightweight [105, 60]. like [99, 4, 475, 105, 252, 441, 386, 250]. limited [145, 306]. limited-magnitude [145]. line [215]. Linear [505, 184, 118, 461, 337, 162, 448, 124, 149, 158, 376, 327, 328, 450, 353, 470, 4, 492, 117, 198, 521, 2, 341, 527, 315, 47, 340, 504, 479, 148, 135, 161, 396, 422, 178, 280, 59, 222, 264, 266, 7, 187, 416, 296, 476, 69, 253, 221, 383, 272, 190, 176, 431, 446, 543, 164, 165, 522, 191, 166, 311, 167, 406]. linearity [542, 161]. Linearized [502, 402]. linkage [510]. local [377, 547]. Localised [192]. locally [428]. location [159, 312]. logarithm [333]. logarithmic [120]. look [115]. looking [62]. Low [458, 278, 395, 202, 415, 506, 188, 153, 387, 183, 283, 540, 38, 200, 515]. low-depth [278]. low-hit-zone [188, 200, 515]. Lower [6, 326, 320, 293, 362]. LPN [141]. LWE [300].

machine [472]. magnitude [509, 145, 403, 488, 382]. Maiorana [504, 350]. Malleability [217]. manipulation [440]. mappings [330]. maps [4, 544, 110]. Masking [100, 500, 251]. mass [172, 430]. matched [159]. Mathematical [295]. mathematics [466]. matrices [22, 71, 483, 29, 25, 395, 31, 214, 105, 21, 30, 32, 24, 319, 33]. Matrix [151, 226, 41, 238, 36]. maximal [170, 513]. Maximum [484, 336, 70, 386, 57, 121, 487]. maximum-length [57, 121]. McFarland [504, 350]. MDS [437, 105, 442, 268, 287]. measure [118, 333]. measures [140, 6, 156]. meet [109, 91]. meet-in-the-middle [109, 91]. Melas [122]. memory [257]. Menezes [358]. Message [227, 39]. method [256, 415, 340, 399, 131, 239]. Methods [295, 116, 200]. metric [281]. metrically [320]. middle [109, 91]. Minimal [543, 522, 463, 396, 476, 120, 383, 36]. Minimum [67, 85, 376, 310, 411, 268, 497, 150]. Missing [205]. mistakes [430]. Modes [87, 147]. Modified [218, 491, 293]. modifying [179]. modular [421]. modulo [184, 142, 110, 138, 191]. MOLS [27]. Monomial [471, 50, 126, 103, 362]. monomials [150]. monotone [261]. most [14]. MUBs [528, 27]. Muller [492, 461, 411, 74, 14, 288]. Multi [516, 250, 208]. Multi-party [516]. multi-structure [250]. multi-user [208]. Multidimensional [47, 91, 253]. multilevel [510]. multinegacirculant [364]. multinomials [34]. Multiple [351, 267, 213, 253]. multiple/multidimensional [253]. multiplication [468, 206]. Multiplicative [89, 356, 277, 420, 35, 542, 97]. multisecret [192]. multisequence [148]. multisequences [2, 230]. multivariate [415, 16]. mutual [211]. Mutually [27, 26, 454, 77]. mystery [316].

Near [322, 79, 201, 229]. near-bent [201]. Near-Optimal [322, 229]. Necessary [116]. negabent [168]. negacirculant [360]. negacyclic [172, 430, 369]. nested [88]. Network [42]. NFSR [57, 121, 373]. NGG [12]. NHZ [518]. Niho [259, 285, 199, 279, 305, 222, 338, 291]. NLFSR [487]. NLFSR-based [487]. NLFSRs [512]. no [391]. no-hit-zone [391]. Non [45, 327, 244, 400, 144, 324, 545, 455, 501, 354, 108, 540, 532, 234, 381]. non-abelian [455]. non-chain [545, 501]. non-commutative [234]. non-cyclic [324]. non-overlapping [244]. non-power-of-two [540, 381]. non-prime [327]. non-randomness [108]. non-resiliency [144]. non-uniform [400]. non-unital [532]. non-weakly [354]. nonbinary [233]. noncommutative [532]. nonexistence [483]. Nonlinear [263, 48, 459, 534, 284, 365, 440, 318]. nonlinearities [17]. Nonlinearity [156, 432, 140, 261, 82, 333, 178, 275, 402, 524, 139, 362, 57, 121]. normal [92]. normality [260]. note [256, 426, 383, 287]. Notes [401]. novel [336]. NTRUSign [51]. null [432]. nulls [40]. number [304, 214, 402]. numbers [16, 154].

obtained [545]. OCDMA [410]. odd

[193, 418, 415, 110, 387, 534, 27, 308, 131, 382]. odd-periodic [387, 382]. often [394]. old [133]. on-line [215]. One [100, 380, 388, 67, 85, 39, 448]. one-dimensional [448]. one-round [39]. One-weight [388]. ones [214]. online [249]. open [112, 133]. operates [345]. operations [147, 87, 97]. optical [410]. Optimal [160, 312, 322, 83, 150, 541, 408, 181, 117, 509, 188, 385, 130, 153, 396, 182, 535, 518, 266, 323, 281, 113, 410, 231, 403, 347, 497, 436, 488, 229, 382, 200, 515]. optimality [211]. Optimization [468]. orbit [324]. order [17, 333, 484, 186, 55, 534, 337, 402, 386, 362, 154, 124, 163, 288, 433]. orders [28]. orientable [334]. oriented [331]. orthogonal [410, 27, 228]. other [100]. output [97]. overheads [281]. overlapping [244]. overview [5].

packings [428]. pair [341, 442, 242]. pairing [101]. pairs [81, 232, 390]. PAM [40]. paper [430]. PAPR [540]. parallel [109, 510]. parallel-cut [109]. parameters [367, 396, 535, 83, 515]. parametrized [226]. Paraunitary [81]. partial [385, 350, 36, 33]. Partially [394]. particular [93]. partitioned [93, 307]. party [516]. PD [462]. Perfect [80, 225, 233, 459, 79, 111, 380, 102, 135, 534, 182]. perfectly [39, 493, 347]. period [148, 254, 318, 488, 229, 487]. Periodic [309, 6, 387, 113, 382, 456]. periods [233, 521, 438]. Permutation [419, 44, 235, 40, 259, 434, 285, 348, 478, 199, 289, 279, 11, 308, 338, 291, 237]. Permutations [451, 504, 546, 432, 517, 353, 459, 349, 520, 465, 136, 297, 284, 179]. perturbed [490]. phase [72, 482]. photonics [458]. PIR [469]. planar [218, 135]. plateaued [429]. PN [34]. POEx [215]. points [517, 346, 171]. Poly [300]. Polynomial [227, 269, 484, 471, 128, 69, 508]. Polynomials [502, 451, 159, 434, 219, 18, 335, 419, 547, 44, 235, 289, 402, 16, 486, 237]. Polyphase [379, 506]. power [415, 478, 520, 30, 27, 540, 508, 229, 519, 381]. practicability [211]. Practical [183]. Preface [526, 157]. Prefix [93]. Preimages [502]. prescribed [304, 214]. presemifields [158]. PRESENT [86]. PRESENT-80 [86]. prime [327, 126, 186, 30, 27, 163, 10, 229, 438]. prime-power [229]. primitive [142, 110, 390, 359, 263, 282, 330]. primitives [100]. privacy [469, 114]. private [114]. Probabilistic [189, 524]. probabilities [102]. probability [153, 253]. problem [273, 26]. problems [133, 422]. product [219, 238, 391]. products [438]. profile [118]. profiles [335]. Progress [21]. progressions [399]. projection [64]. Properties [392, 429, 255, 42, 544, 143, 137, 153, 491, 96, 88, 49, 421]. property [153, 314]. proposal [106]. protected [471]. protocols [114]. provable [257]. proven [49]. proving [256]. pseudo [523]. pseudo-ultrametric [523]. Pseudorandom [539, 16]. pseudorandomness [258]. public [301]. public-key [301]. punctured [411].

QAM [81, 381]. QC [405]. QC-LDPC [405]. Quadratic [224, 37, 134, 416, 486, 185, 167]. Quadruple [366]. Quantum [292, 480, 499, 498, 545, 501, 530, 238, 477, 334, 268]. quartic [286]. Quasi [159, 135, 322, 477, 125, 357, 453, 317, 228, 174]. Quasi-Complementary [322]. Quasi-cyclic [159, 477, 125, 357, 453, 228, 174]. Quasi-perfect [135]. quasi-twisted [317]. Quaternary [64, 541, 408, 184, 117, 521, 153, 283, 220, 38, 382, 406]. quaternions [80, 225]. Quiescent [458]. quotients [43, 456].

R [329]. R-2 [329]. Rabbit [59]. radii [328]. radius [288]. ramp [73]. random [156]. randomisation [253]. randomness [525, 108, 329]. rank [404, 281]. rank-metric [281]. rate [249]. rate-1 [249]. Rational [523, 478, 402, 40]. RC4 [316, 232, 251, 11, 254, 108]. RC4A [108]. readout [458]. real [454, 528]. realistic [56]. realization [36]. REALLY [31]. recovery [12]. recurring [2, 7]. recursive [263]. reduced [8]. Reed [492, 536, 461, 411, 74, 14, 268, 288]. Reed-Muller [74, 14]. registers [226, 230]. Regular [428, 343, 392, 162, 320, 354, 522]. regularity [355]. Related [116, 86, 18, 133, 530]. related-key [86]. relations [363]. Relative [335, 225, 35, 332, 169, 447]. remainder [125]. remarks [32]. repairable [428]. Repeated [529, 104, 439, 242]. Repeated-root [104, 439, 242]. representation [184, 456]. representations [151, 309]. residue [224, 185]. Resilience [75]. resiliency [144, 500]. Resilient [500, 100, 441]. resist [96]. resistance [86]. resistant [46, 61]. Resolvable [528]. respect [262, 421]. restricted [275]. Results [313, 483, 464, 422, 114, 121, 338, 237]. retrieval [114]. reverse [99]. reverse-engineering [99]. Reversed [137]. reversible [474]. Revisiting [90, 232]. ring [377, 545, 474, 224, 501, 183, 453, 317, 532, 234]. rings [366, 409, 547]. rise [353]. robust [210, 107, 331]. Root [425, 377, 529, 104, 439, 242]. Root-Hadamard [425]. rotation [50, 66, 126]. round [129, 39]. RSBFs [96]. RSSBs [313]. Rudin [386]. rule [68].

s [100, 48, 95, 475, 137, 542, 313, 178, 441, 274, 298, 89]. S-boxes [48, 95, 475, 137, 542, 313, 178, 441, 274, 298, 89]. scalable [68]. scalar [468]. SCARE [99]. scenarios [56, 211]. scheme [207, 469, 152, 51]. schemes [496, 392, 84, 281, 209, 73, 194, 65]. search [232, 405]. Searchable [213]. Second [74, 186, 169, 534, 402, 14, 362, 288, 433]. second-order [534, 402, 362]. Secondary [70]. Secret [299, 496, 84, 194, 210, 39, 11, 65]. sections [203]. Secure [227, 215, 249, 39, 281]. Security [152, 208, 256, 301, 257, 331, 472]. Self [444, 366, 474, 536, 409, 310, 485, 547, 417, 352, 317, 127, 228, 360, 220, 287, 457]. Self-dual [444, 366, 474, 536, 409, 310, 547, 417, 352, 317, 127, 228, 360, 220, 287, 457]. self-orthogonal [228]. semifields [34]. sensing [395, 319]. separable [181, 336, 236]. Separation [464, 332]. Sequence [322, 81, 223, 180, 153, 186, 397, 518, 241, 540, 399, 293, 386, 229, 200, 314, 391, 381, 515]. Sequences [76, 201, 225, 197, 505, 408, 233, 219, 184, 142, 117, 198, 521, 509, 2, 384, 143, 506, 188, 385, 79, 118, 378, 110, 153, 6, 380, 387, 525, 337, 425, 7, 187, 258, 539, 283, 40, 427, 365, 69, 231, 403, 246, 329, 263, 523, 36, 124, 170, 282, 330, 393, 318, 488, 10, 438, 38, 382, 379, 456, 406, 516]. server [469]. servers [213]. Set [322, 110, 417, 28]. sets [225, 408, 342, 404, 93, 159, 355, 462, 3, 72, 35, 188, 145, 518, 540, 229, 200, 314, 391, 381, 515]. setting [208]. Settling [316]. Several [199, 280, 352, 476, 291, 433, 431, 325]. SHA [8]. SHA-2 [8]. Shapiro [386].

Shapiro-like [386]. sharing [299, 496, 84, 210, 192, 39, 399, 298, 194, 65]. shift [226, 230]. Shifted [429]. shortest [486]. side [255, 458, 472]. side-channel [255, 472]. Sidelnikov [197]. signal [72]. signals [40]. signature [51, 189]. signatures [208, 120]. significant [105, 333, 139]. similarities [63]. simple [78, 377, 73, 120, 212, 49]. simplex [442, 492]. single [469, 306]. single-server [469]. Singly [310]. six [460, 405]. six-valued [460]. size [306, 320]. sizes [80, 67, 85, 390]. Skew [453, 125]. SKINNY [475]. SKINNY-like [475]. Small [278, 3, 95, 414, 252, 421]. small-state [252]. smallest [405]. Solomon [536, 268]. Solving [412, 141]. Some [31, 247, 116, 491, 182, 283, 407, 32, 268, 163, 533, 158, 505, 172, 426, 517, 434, 430, 459, 520, 270, 245, 547, 401, 7, 167, 211]. space [309, 410]. sparse [528, 339]. Special [466, 344, 271, 295, 413, 489, 157, 76, 204, 378, 526, 159, 424, 20, 317, 132]. Spectra [134, 460, 534, 325]. spectral [40, 398]. spectrum [148]. spectrums [520]. splitting [23, 102, 182]. sporadic [120]. spread [350]. spreads [158]. square [256, 110, 454]. square-free [110]. squares [27]. SRAM [458]. SRs [523]. standard [209]. Stanica [112]. State [58, 252, 309, 254]. state-space [309]. Statistical [255, 216, 250, 143, 251, 329, 473]. stealthy [473]. Steiner [3]. step [8]. stopping [3]. stream [68, 56, 173, 63, 252, 257, 5, 12, 13, 55, 9, 83, 96, 60, 294, 87, 189, 88, 49, 58, 487]. strikes [415]. Strong [455, 349]. Strongly [354, 236]. structural [63, 13]. structure [269, 42, 250, 357, 373]. structures [504, 39, 416, 194]. study [545]. subclass [17]. subcodes [461]. subfield [431, 446, 443]. subfields [179]. submultiples [40]. suboptimal [78]. subsequences [484]. subset [320]. subsets [390]. subspace [503, 510, 449]. Success [253]. Sufficient [116]. sum [404, 478, 484, 246]. sums [426, 243, 397, 448, 490, 221, 196, 343, 303]. supplementary [408]. support [511]. supporting [87]. surfaces [334]. Survey [537, 54, 101, 305, 538]. Suzuki [120]. swapping [517]. switching [131]. symbol [442, 40, 123, 242]. symbol-pair [442, 242]. symbols [365]. symmetric [50, 356, 66, 126, 509, 313, 107, 213]. synchronizable [480]. synthesis [230]. Synthetic [59]. systems [3, 173, 290, 410].

tables [62]. Tangent [544]. Tangent-Chebyshev [544]. techniques [527]. terms [343]. ternary [380, 497, 393, 519]. Teske [358]. Test [227]. tests [525, 524, 329]. th [333, 386]. their [218, 429, 132, 344, 271, 413, 489, 307, 418, 462, 75, 224, 76, 378, 21, 64, 428, 460, 479, 529, 511, 136, 104, 24, 154, 446, 497, 513, 166]. theorem [125]. theory [526, 90, 255, 20, 245, 270]. Third [17]. Third-order [17]. Three [395, 482, 332, 387, 222, 264, 410, 195, 176, 165, 166]. three-dimensional [410]. Three-phase [482]. three-weight [332, 176, 166]. Threshold [95, 471, 481, 73]. tight [392]. Tighter [518]. time [72, 257, 410]. time-memory-data [257]. time-phase [72]. too [212]. tool [256]. Trace [456, 304, 184, 436]. tradeoff [257]. transform [262, 103]. transformation [512]. transforms [146, 425, 351]. transition [254]. translators [353]. transmission [39]. treatment [73]. trellis [447]. trends [204, 537]. triads [482]. trick [205]. trinomials [259, 285, 348, 199, 235, 279, 308, 338, 291]. triple [3]. Trivium [58]. Trojan [473]. Truncated [267]. truth [62]. tuple [241]. tweakable [249]. twin [130, 30, 438, 152]. Twin-Beth [152]. twin-free [130]. twisted [423, 317]. Two [323, 195, 436, 443, 165, 229, 381, 219, 177, 346, 186, 306, 161, 222, 266, 361, 317, 540, 388, 370, 221, 508, 154, 190, 488, 438, 311]. two-error-correcting [361]. two-prime [186]. two-weight [266, 388, 221, 311]. type [494].

ultrametric [523]. Unbalanced [42].

Unbiased [25, 26, 454, 27, 77]. unbounded [80]. unequal [390]. Unicyclic [349]. uniform [240, 432, 400, 435, 179, 298]. uniformities [537]. uniformity [434, 500, 415, 542, 370, 490, 139, 519]. Unit [72]. unital [532]. Universal [281, 336]. Upper [356, 290, 267]. user [208, 114]. user-private [114]. users [114]. using [141, 255, 198, 249, 302, 87, 77]. UWB [384].

v1 [129]. value [117]. valued [460]. values [197]. variable [277]. variants [141]. variations [205]. varieties [203]. variety [286]. various [140]. varying [36]. vector [450, 87]. Vectorial [418, 82, 151, 263]. version [98]. vertex [67]. via [527, 514, 440, 436, 131, 433]. violating [476]. VMPC [108]. Volume [1].

Walsh [92, 460, 103, 146]. watermark [516]. watermarks [111]. wave [410]. wave-length [410]. wave-length/space/time [410]. weak [300]. weakly [162, 354, 522]. weakness [13]. weighing [22, 71, 483]. weight [450, 332, 310, 177, 335, 479, 74, 169, 264, 266, 442, 447, 14, 195, 388, 221, 190, 176, 166, 311, 167]. Weights [50, 376, 62, 280, 401, 222, 162, 164, 443, 165]. weightwise [493, 347]. Weil [490]. Weng [358]. WG [83, 60, 294, 75]. WG-16 [294]. WG-7 [60, 75]. WG-8 [294]. Whiteman [186, 480]. wireless [173]. wise [435, 447]. within [372, 371]. without [62, 227, 504].

XCB [115]. XOR [123].

Zeckendorf [484]. zero [404, 534, 239, 379]. zero-difference [239]. zero-sum [404]. zeros [402]. zone [188, 379, 200, 391, 515].

References

Carlet:2009:ECC

[1] Claude Carlet. Editorial: Cryptography and Communications, volume 1, issue 1. Cryptography and Communications, 1(1):1–2, April 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s12095-009-0010-6.pdf.

Fu:2009:JLC

[2] Fang-Wei Fu, Harald Niederreiter, and Ferruh Özbudak. Joint linear complexity of multisequences consisting of linear recurring sequences. Cryptography and Communications, 1(1): 3–29, April 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-007-0001-4.

Colbourn:2009:SSS

[3] Charles J. Colbourn and Yuichiro Fujiwara. Small stopping sets in Steiner triple systems. Cryptography and Communications, 1(1):31–46, April 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-008-0002-
y.

Daemen:2009:NCL

[4] Joan Daemen and Vincent Rijmen. New criteria for linear maps in AES-like ciphers. Cryptography and Communications, 1(1):47–69, April 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-008-0003-x.

Hell:2009:ODA

[5] Martin Hell, Thomas Johansson, and Lennart Brynielsson. An overview of distinguishing attacks on stream ciphers. Cryptography and Communications, 1(1):71–94, April 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-008-0006-7.

Kavuluru:2009:LBE

[6] Ramakanth Kavuluru and Andrew Klapper. Lower bounds on error complexity measures for periodic LFSR and FCSR sequences. Cryptography and Communications, 1(1): 95–116, April 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-008-0004-9.

Meidl:2009:HDL

[7] Wilfried Meidl. How to determine linear complexity and k-error linear complexity in some classes of linear recurring sequences. Cryptography and Communications, 1(1):117–133, April 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-008-0007-6.

Sanadhya:2009:CAR

[8] Somitra Kumar Sanadhya and Palash Sarkar. A combinatorial analysis of recent attacks on step reduced SHA-2 family. Cryptography and Communications, 1(2):135–173, September 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-009-0011-5.

Kolokotronis:2009:FDF

[9] Nicholas Kolokotronis, Konstantinos Limniotis, and Nicholas Kalouptsidis. Factorization of determinants over finite fields and application in stream ciphers. Cryptography and Communications, 1 (2):175–205, September 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-008-0005-8.

Xu:2009:ASP

[10] Hong Xu, Wen-Feng Qi, and Yong-Hui Zheng. Autocorrelations of l-sequences with prime connection integer. Cryptography and Communications, 1(2): 207–223, September 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-008-0008-5.

Paul:2009:BPK

[11] Goutam Paul and Subhamoy Maitra. On biases of permutation and keystream bytes of RC4 towards the secret key. Cryptography and Communications, 1 (2):225–268, September 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-008-0009-4.

Kircanski:2009:NDK

[12] Aleksandar Kircanski, Rabeah Al-Zaidy, and Amr M. Youssef. A new distinguishing and key recovery attack on NGG stream cipher. Cryptography and Communications, 1(2): 269–282, September 2009. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-009-0012-4.

Kircanski:2010:SWG

[13] Aleksandar Kircanski and Amr M. Youssef. On the structural weakness of the GGHN stream cipher. Cryptography and Communications, 2 (1):1–17, April 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-009-0013-3.

Rolland:2010:SWG

[14] Robert Rolland. The second weight of generalized Reed-Muller codes in most cases. Cryptography and Communications, 2(1):19–40, April 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-009-0014-2.

Pelto:2010:NBI

[15] Mikko Pelto. New bounds for (r, ≤ 2)- identifying codes in the infinite king grid. Cryptography and Communications, 2(1):41–47, April 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-009-0015-1.

Ostafe:2010:PNH

[16] Alina Ostafe and Igor E. Shparlinski. Pseudorandom numbers and hash functions from iterations of multivariate polynomials. Cryptography and Communications, 2(1):49–67, April 2010. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-009-0016-
0.

Gode:2010:TON

[17] Ruchi Gode and Sugata Gangopadhyay. Third-order nonlinearities of a subclass of Kasami functions. Cryptography and Communications, 2(1): 69–83, April 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-009-0017-z.

Helleseth:2010:BXR

[18] Tor Helleseth and Alexander Kholosha. x^{2^{l}+1} + x + a and related affine polynomials over GF(2^{k}). Cryptography and Communications, 2(1): 85–109, April 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-009-0018-y.

Canfield:2010:AEC

[19] E. Rodney Canfield, Zhicheng Gao, Catherine Greenhill, Brendan D. McKay, and Robert W. Robinson. Asymptotic enumeration of correlationimmune Boolean functions. Cryptography and Communications, 2(1):

111–126, April 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-010-0019-x.

Flannery:2010:GES

[20] Dane L. Flannery and Kathryn J. Horadam. Guest editorial for the special issue on design theory. Cryptography and Communications, 2(2): 127–128, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s12095-010-0035-x.pdf.

Horadam:2010:HMT

[21] K. J. Horadam. Hadamard matrices and their applications: Progress 2007–2010. Cryptography and Communications, 2 (2):129–154, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-010-0032-0.

Arasu:2010:CWM

[22] Krishnasamy Thiru Arasu and Alex J. Gutman. Circulant weighing matrices. Cryptography and Communications, 2 (2):155–171, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-010-0025-z.

Huber:2010:CBC

[23] Michael Huber. Combinatorial bounds and characterizations of splitting authentication codes. Cryptography and Communications, 2(2):173–185, September 2010. CODEN ????

ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-010-0020-4.

Szollosi:2010:ECH

[24] Ferenc Szöllősi. Exotic complex Hadamard matrices and their equivalence. Cryptography and Communications, 2(2):187–198, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-010-0021-
3.

Best:2010:UCH

[25] Darcy Best and Hadi Kharaghani. Unbiased complex Hadamard matrices and bases. Cryptography and Communications, 2(2):199–209, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-010-0029-
8.

Jaming:2010:PMU

[26] Philippe Jaming, Máté Matolcsi, and Péter Móra. The problem of mutually unbiased bases in dimension 6. Cryptography and Communications, 2 (2):211–220, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-010-0023-1.

Rao:2010:MOL

[27] Asha Rao, Diane Donovan, and Joanne L. Hall. Mutually orthogonal Latin squares and mutually unbiased bases in dimensions of odd prime power MOLS and MUBs in odd prime power dimensions. Cryptography and Communications, 2(2): 221–231, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-010-0027-x.

deLauney:2010:DSK

[28] Warwick de Launey and Daniel M. Gordon. On the density of the set of known Hadamard orders. Cryptography and Communications, 2(2): 233–246, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-010-0028-9.

Armario:2010:ICC

[29] José Andrés Armario. On an inequivalence criterion for cocyclic Hadamard matrices. Cryptography and Communications, 2(2):247–259, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s12095-010-0024-
                                        0.

OCathain:2010:TPP

[30] Padraig Ó Catháin and Richard M. Stafford. On twin prime power Hadamard matrices. Cryptography and Communications, 2(2):261–269, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-010-0030-2.

Dillon:2010:SRB

[31] J. F. Dillon. Some REALLY beautiful Hadamard matrices. Cryptography and Communications, 2(2): 271–292, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-010-0031-1.

Seberry:2010:SRH

[32] Jennifer Seberry and Marilena Mitrouli. Some remarks on Hadamard matrices. Cryptography and Communications, 2 (2):293–306, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-010-0036-9.

deLauney:2010:FAA

[33] Warwick de Launey and David A. Levin. A Fourier-analytic approach to counting partial Hadamard matrices. Cryptography and Communications, 2(2): 307–334, September 2010. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-010-0033-z.

Budaghyan:2011:NCS

[34] Lilya Budaghyan and Tor Helleseth. New commutative semifields defined by new PN multinomials. Cryptography and Communications, 3(1): 1–16, March 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-010-0022-2.

Farmer:2011:ECM

[35] D. G. Farmer and K. J. Horadam. Equivalence classes of multiplicative central (p^{n}, p^{n}, p^{n}, 1)-relative difference sets. Cryptography and Communications, 3(1):17–28, March 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-010-0026-y.

Wang:2011:LBM

[36] Li-Ping Wang. A lattice-based minimal partial realization algorithm for matrix sequences of varying length. Cryptography and Communications, 3 (1):29–42, March 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-010-0037-8.

Bracken:2011:FMQ

[37] Carl Bracken, Eimear Byrne, Nadya Markin, and Gary McGuire. A few more quadratic APN functions. Cryptography and Communications, 3 (1):43–53, March 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-010-0038-7.

Yang:2011:CQS

[38] Zheng Yang and Pinhui Ke. Construction of quaternary sequences of length pq with low autocorrelation. Cryptography and Communications, 3(2):55–64, June 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-010-0034-y.

Martin:2011:EDS

[39] Keith M. Martin, Maura B. Paterson, and Douglas R. Stinson. Error decodable secret sharing and oneround perfectly secure message transmission for general adversary structures. Cryptography and Communications, 3(2):65–86, June 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-010-0039-6.

Ouahada:2011:PSC

[40] Khmaies Ouahada, Theo G. Swart, and Hendrik C. Ferreira. Permutation sequences and coded PAM signals with spectral nulls at rational submultiples of the symbol frequency. Cryptography and Communications, 3(2):87–108, June 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-010-0040-0.

Arnault:2011:MAF

[41] François Arnault, Thierry P. Berger, and Benjamin Pousse. A matrix approach for FCSR automata. Cryptography and Communications, 3(2): 109–139, June 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-010-0041-z.

Choy:2011:CPA

[42] Jiali Choy, Guanhan Chew, Khoongming Khoo, and Huihui Yap. Cryptographic properties and application of a Generalized Unbalanced Feistel Network structure. Cryptography and Communications, 3(3):141– 164, September 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-011-0042-6.

Aly:2011:BFD

[43] Hassan Aly and Arne Winterhof. Boolean functions derived from Fermat quotients. Cryptography and Communications, 3(3):165–174, September 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s12095-011-0043-
                                        5.

Li:2011:PPE

[44] Yongqiang Li and Mingsheng Wang. Permutation polynomials EA-equivalent to the inverse function over GF(2^{n}). Cryptography and Communications, 3 (3):175–186, September 2011. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-011-0045-3.

Poinsot:2012:NAB

[45] Laurent Poinsot. Non Abelian bent functions. Cryptography and Communications, 4(1):1–23, March 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-011-0058-y.

Pasalic:2012:DBF

[46] Enes Pasalic. A design of Boolean functions resistant to (fast) algebraic cryptanalysis with efficient implementation. Cryptography and Communications, 4(1):25–45, March 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-011-0057-z. See com-

ments [61].

Hermelin:2012:MLD

[47] Miia Hermelin and Kaisa Nyberg. Multidimensional linear distinguishing attacks and Boolean functions. Cryptography and Communications, 4 (1):47–64, March 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-011-0053-3.

Beelen:2012:NCH

[48] Peter Beelen and Gregor Leander. A new construction of highly nonlinear Sboxes. Cryptography and Communications, 4(1):65–77, March 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-011-0052-4.

Si:2012:SSC

[49] Wenpei Si and Cunsheng Ding. A simple stream cipher with proven properties. Cryptography and Communications, 4(2):79–104, June 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-011-0059-x.

Bileschi:2012:WBC

[50] Maxwell L. Bileschi, Thomas W. Cusick, and Daniel Padgett. Weights of Boolean cubic monomial rotation symmetric functions. Cryptography and Communications, 4(2): 105–130, June 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-011-0060-4.

Kamal:2012:FAN

[51] Abdel Alim Kamal and Amr M. Youssef. Fault analysis of the NTRUSign digital signature scheme. Cryptography and Communications, 4 (2):131–144, June 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-011-0061-3.

Maitra:2012:GE

[52] Subhamoy Maitra and Palash Sarkar. Guest editorial. Cryptography and Communications, 4(3–4):145–146, December 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s12095-
                                        012-0073-7.pdf.

Canteaut:2012:CAC

[53] Anne Canteaut and María Naya-Plasencia. Correlation attacks on combination generators. Cryptography and Communications, 4(3–4): 147–171, December 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-012-0069-3.

Aagren:2012:SFC

[54] Martin Ågren, Carl Löndahl, Martin Hell, and Thomas Johansson. A survey on fast correlation attacks. Cryptography and Communications, 4(3–4): 173–202, December 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-012-0062-x.

Knellwolf:2012:HOD

[55] Simon Knellwolf and Willi Meier. High order differential attacks on stream ciphers. Cryptography and Communications, 4(3–4):203–215, December 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-012-0071-
9.

Dinur:2012:ACA

[56] Itai Dinur and Adi Shamir. Applying cube attacks to stream ciphers in realistic scenarios. Cryptography and Communications, 4(3–4): 217–232, December 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-012-0068-4.

Turan:2012:NML

[57] Meltem Sönmez Turan. On the nonlinearity of maximum-length NFSR feedbacks. Cryptography and Communications, 4(3–4):233–243, December 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-012-0067-
5.

Simpson:2012:SCI

[58] Leonie Simpson and Serdar Boztas. State cycles, initialization and the Trivium stream cipher. Cryptography and Communications, 4(3–4): 245–258, December 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-012-0066-6.

Lu:2012:SLA

[59] Yi Lu, Serge Vaudenay, and Willi Meier. Synthetic linear analysis with applications to CubeHash and Rabbit. Cryptography and Communications, 4(3–4): 259–276, December 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-012-0064-8.

Orumiehchiha:2012:CWL

[60] Mohammad Ali Orumiehchiha, Josef Pieprzyk, and Ron Steinfeld. Cryptanalysis of WG-7: a lightweight stream cipher. Cryptography and Communications, 4(3–4):277–285, December 2012. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-012-0070-
x.

Wang:2013:CDB

[61] Wenhao Wang, Meicheng Liu, and Yin Zhang. Comments on “A design of Boolean functions resistant to (fast) algebraic cryptanalysis with efficient implementation”. Cryptography and Communications, 5(1):1–6, March 2013. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-012-0063-

9. See [46].

Cusick:2013:FHW

[62] Thomas W. Cusick. Finding Hamming weights without looking at truth tables. Cryptography and Communications, 5(1):7–18, March 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-012-0072-8.

Gupta:2013:DIA

[63] Sourav Sen Gupta, Anupam Chattopadhyay, and Ayesha Khalid. Designing integrated accelerator for stream ciphers with structural similarities. Cryptography and Communications, 5 (1):19–47, March 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-012-0074-6.

Jadda:2013:QCB

[64] Zoubida Jadda, Patrice Parraud, and Soukayna Qarboua. Quaternary cryptographic bent functions and their binary projection. Cryptography and Communications, 5(1):49–65, March 2013. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-012-0077-
3.

delaCruz:2013:CIS

[65] Romar dela Cruz and Huaxiong Wang. Cheating-immune secret sharing schemes from codes and cumulative arrays. Cryptography and Communications, 5(1):67–83, March 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-012-0076-4.

Brown:2013:ECC

[66] Alyssa Brown and Thomas W. Cusick. Equivalence classes for cubic rotation symmetric functions. Cryptography and Communications, 5(2): 85–118, June 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-012-0075-5.

Charon:2013:MSI

[67] Irène Charon, Iiro Honkala, Olivier Hudry, and Antoine Lobstein. Minimum sizes of identifying codes in graphs differing by one vertex. Cryptography and Communications, 5(2): 119–136, June 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-012-0078-2.

Das:2013:CNS

[68] Sourav Das and Dipanwita RoyChowdhury. CAR30: a new scalable stream cipher with rule 30. Cryptography and Communications, 5(2): 137–162, June 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-012-0079-1.

Salagean:2013:CLC

[69] Ana Sălăgean, Alex J. Burrage, and Raphael C.-W. Phan. Computing the linear complexity for sequences with characteristic polynomial f^{v}. Cryptography and Communications, 5 (2):163–177, June 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0080-3.

Limniotis:2013:SCB

[70] Konstantinos Limniotis, Nicholas Kolokotronis, and Nicholas Kalouptsidis. Secondary constructions of Boolean functions with maximum algebraic immunity. Cryptography and Communications, 5(3):179–199, September 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-013-0081-
2.

Arasu:2013:BWM

[71] K. T. Arasu, Simone Severini, and Edmund Velten. Block weighing matrices. Cryptography and Communications, 5 (3):201–207, September 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0083-0.

Ding:2013:UTP

[72] Cunsheng Ding, Keqin Feng, Rongquan Feng, Maosheng Xiong, and Aixian Zhang. Unit time-phase signal sets: Bounds and constructions. Cryptography and Communications, 5(3): 209–227, September 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0085-y.

Paterson:2013:SCT

[73] Maura B. Paterson and Douglas R. Stinson. A simple combinatorial treatment of constructions and threshold gaps of ramp schemes. Cryptography and Communications, 5(4):

229–240, December 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-013-0082-1.

Leducq:2013:SWC

[74] Elodie Leducq. Second weight codewords of generalized Reed-Muller codes. Cryptography and Communications, 5 (4):241–276, December 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-013-0084-z.

Gong:2013:RDA

[75] Guang Gong, Mark Aagaard, and Xinxin Fan. Resilience to distinguishing attacks on WG-7 cipher and their generalizations. Cryptography and Communications, 5(4): 277–289, December 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-013-0089-7.

Helleseth:2014:SIE

[76] Tor Helleseth and Jonathan Jedwab. Special issue editorial: Sequences and their applications. Cryptography and Communications, 6(1):1–2, March 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s12095-014-0098-1.pdf.

Wu:2014:CCU

[77] Gaofei Wu and Matthew Geoffrey Parker. A complementary construction using mutually unbiased bases. Cryptography and Communications, 6 (1):3–25, March 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0095-9.

Bajic:2014:SSC

[78] Dragana Bajic and Tatjana Loncar-Turukalo. A simple suboptimal construction of cross-bifix-free codes. Cryptography and Communications, 6(1): 27–37, March 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0088-8.

Hariharan:2014:NNP

[79] Rema Hariharan. New near perfect sequences of even lengths. Cryptography and Communications, 6(1): 39–46, March 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0093-y.

Acevedo:2014:PAU

[80] Santiago Barrera Acevedo and Nathan Jolly. Perfect arrays of unbounded sizes over the basic quaternions. Cryptography and Communications, 6(1): 47–57, March 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0086-x.

Budisin:2014:PGC

[81] S. Z. Budišin and P. Spasojević. Paraunitary generation/correlation of QAM complementary sequence pairs. Cryptography and Communications, 6 (1):59–102, March 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0087-9.

Dib:2014:ANV

[82] Stéphanie Dib. Asymptotic nonlinearity of vectorial Boolean functions. Cryptography and Communications, 6 (2):103–115, June 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0090-1.

Mandal:2014:OPW

[83] Kalikinkar Mandal, Guang Gong, Xinxin Fan, and Mark Aagaard. Optimal parameters for the WG stream cipher family. Cryptography and Communications, 6(2):117–135, June 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-013-0091-
0.

Gao:2014:SSS

[84] Ying Gao and Romar dela Cruz. Secret sharing schemes based on graphical codes. Cryptography and Communications, 6(2):137–155, June 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0092-z.

Charon:2014:MSI

[85] Irène Charon, Iiro Honkala, Olivier Hudry, and Antoine Lobstein. Minimum sizes of identifying codes in graphs differing by one edge. Cryptography and Communications, 6(2):

157–170, June 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0094-x.

Emami:2014:RPA

[86] Sareh Emami, San Ling, Ivica Nikolić, Josef Pieprzyk, and Huaxiong Wang. The resistance of PRESENT-80 against related-key differential attacks. Cryptography and Communications, 6(3): 171–187, September 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0096-8.

Sarkar:2014:MOE

[87] Palash Sarkar. Modes of operations for encryption and authentication using stream ciphers supporting an initialisation vector. Cryptography and Communications, 6(3): 189–231, September 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-013-0097-7.

Shan:2014:CPN

[88] Jinyong Shan, Lei Hu, and Xiangyong Zeng. Cryptographic properties of nested functions and algebraic immunity of the Boolean function in Hitag2 stream cipher. Cryptography and Communications, 6(3): 233–254, September 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-014-0099-0.

Zajac:2014:MCB

[89] Pavol Zajac and Matúš Jókay. Multiplicative complexity of bijective 4 × 4 S-boxes. Cryptography and Communications, 6(3):255–277, September 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s12095-014-0100-
                                        y.

Bay:2014:RIA

[90] AslıBay, Atefeh Mashatan, and Serge Vaudenay. Revisiting iterated attacks in the context of decorrelation theory. Cryptography and Communications, 6 (4):279–311, December 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-014-0101-x.

Zhu:2014:MMM

[91] Bo Zhu and Guang Gong. Multidimensional meet-in-the-middle attack and its applications to KATAN32/48/64. Cryptography and Communications, 6 (4):313–333, December 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-014-0102-9.

Gong:2014:CWC

[92] Xinxin Gong, Bin Zhang, Wenling Wu, and Dengguo Feng. Computing Walsh coefficients from the algebraic normal form of a Boolean function. Cryptography and Communications, 6(4): 335–358, December 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-014-0103-8.

Bernini:2014:PPG

[93] Antonio Bernini, Stefano Bilotta, Renzo Pinzani, Ahmad Sabri, and Vincent Vajnovszki. Prefix partitioned gray codes for particular cross-bifix-free sets. Cryptography and Communications, 6 (4):359–369, December 2014. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-014-0105-6.

Carlet:2015:GE

[94] Claude Carlet and Pierre-Alain Fouque. Guest editorial. Cryptography and Communications, 7(1):1–2, March 2015. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s12095-014-0115-4.pdf.

Bilgin:2015:TIS

[95] Begül Bilgin, Svetla Nikova, Ventzislav Nikov, Vincent Rijmen, Natalia Tokareva, and Valeriya Vitkup. Threshold implementations of small S-boxes. Cryptography and Communications, 7(1): 3–33, March 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-014-0104-7.

Mazumdar:2015:CRI

[96] Bodhisatwa Mazumdar, Debdeep Mukhopadhyay, and Indranil Sengupta. Construction of RSBFs with improved cryptographic properties to resist differential fault attack on grain family of stream ciphers. Cryptography and Communications, 7(1):35–69, March 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-014-0108-3.

Tunstall:2015:DIB

[97] Michael Tunstall and Marc Joye. The distributions of individual bits in the output of multiplicative operations. Cryptography and Communications, 7 (1):71–90, March 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-014-0110-9.

Bauer:2015:HCC

[98] Aurélie Bauer, Eliane Jaulmes, Emmanuel Prouff, Jean-René Reinhard, and Justine Wild. Horizontal collision correlation attack on elliptic curves—extended version. Cryptography and Communications, 7(1):91– 119, March 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-014-0111-8.

Clavier:2015:CRE

[99] Christophe Clavier, Quentin Isorez, Damien Marion, and Antoine Wurcker. Complete reverse-engineering of AESlike block ciphers by SCARE and FIRE attacks. Cryptography and Communications, 7(1):121–162, March 2015. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s12095-014-0112-

Belaid:2015:MLR

[100] Sonia Belaïd, Vincent Grosso, and François-Xavier Standaert. Masking and leakage-resilient primitives: One, the other(s) or both? Cryptography and Communications, 7(1): 163–184, March 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-014-0113-6.

ElMrabet:2015:SFA

[101] Nadia El Mrabet, Jacques J. A.

Fournier, Louis Goubin, and Ronan Lashermes. A survey of fault attacks in pairing based cryptography. Cryptography and Communications, 7 (1):185–205, March 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-014-0114-5.

Li:2015:CFP

[102] Mingchao Li, Miao Liang, and Beiliang Du. A construction of t-fold perfect splitting authentication codes with equal deception probabilities. Cryptography and Communications, 7 (2):207–215, June 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-014-0107-4.

Li:2015:WTC

[103] Chengju Li and Qin Yue. The Walsh transform of a class of monomial functions and cyclic codes. Cryptography and Communications, 7(2): 217–228, June 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-014-0109-2.

Sharma:2015:RRC

[104] Anuradha Sharma. Repeated-root constacyclic codes of length ℓ^{t}p^{s} and their dual codes. Cryptography and Communications, 7(2):229–255, June 2015. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s12095-014-0106-
                                        5.

Gupta:2015:CSM

[105] Kishan Chand Gupta and Indranil Ghosh Ray. Cryptographically significant MDS matrices based on circulant and circulant-like matrices for lightweight applications. Cryptography and Communications, 7(2):257–287, June 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s12095-014-0116-
                                        3.

Luo:2015:ADL

[106] Yiyuan Luo, Xuejia Lai, and Tiejun Jia. Attacks on a double length blockcipher-based hash proposal. Cryptography and Communications, 7(3): 289–295, September 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-014-0117-2.

Li:2015:CRI

[107] Yuan Li. Characterization of robust immune symmetric boolean functions. Cryptography and Communications, 7 (3):297–315, September 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-014-0120-7.

Sarkar:2015:FNR

[108] Santanu Sarkar. Further non-randomness in RC4, RC4A and VMPC. Cryptography and Communications, 7(3): 317–330, September 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-014-0119-0.

Nikolic:2015:PCM

[109] Ivica Nikolić, Lei Wang, and Shuang Wu. The parallel-cut meet-in-the-middle attack. Cryptography and Communications, 7(3):331–345, September 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-014-0118-
1.

Hu:2015:ICM

[110] Zhi Hu and Lin Wang. Injectivity of compressing maps on the set of primitive sequences modulo square-free odd integers. Cryptography and Communications, 7(4):347– 361, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-014-0121-6.

Jolly:2015:AAA

[111] Nathan Jolly. An algebra of arrays and almost perfect watermarks. Cryptography and Communications, 7(4): 363–377, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-015-0123-z.

Castro:2015:DAO

[112] Francis N. Castro, Oscar E. González, and Luis A. Medina. A divisibility approach to the open boundary cases of Cusick–Li–Stănică’s conjecture. Cryptography and Communications, 7(4): 379–402, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-015-0124-y.

Ortiz-Ubarri:2015:NFA

[113] José Ortiz-Ubarri. New families of asymptotically optimal doubly periodic arrays with ideal correlation constraints. Cryptography and Communications, 7 (4):403–414, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-015-0122-0.

Swanson:2015:ERP

[114] Colleen M. Swanson and Douglas R. Stinson. Extended results on privacy against coalitions of users in user-private information retrieval protocols. Cryptography and Communications, 7(4): 415–437, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-015-0125-x.

Chakraborty:2015:ALX

[115] Debrup Chakraborty, Vicente Hernandez-Jimenez, and Palash Sarkar. Another look at XCB. Cryptography and Communications, 7(4):439– 468, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0127-8.

Hodzic:2015:GBF

[116] S. Hodžić and E. Pasalic. Generalized bent functions — some general construction methods and related necessary and sufficient conditions. Cryptography and Communications, 7(4): 469–483, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0126-9.

Edemskiy:2015:LCB

[117] Vladimir Edemskiy and Andrey Ivanov.

The linear complexity of balanced quaternary sequences with optimal autocorrelation value. Cryptography and Communications, 7(4):485– 496, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0130-0.

He:2015:LCP

[118] Jing Jane He, Daniel Panario, Qiang Wang, and Arne Winterhof. Linear complexity profile and correlation measure of interleaved sequences. Cryptography and Communications, 7(4): 497–508, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0131-z.

Guo:2015:IDE

[119] Chun Guo and Dongdai Lin. Improved domain extender for the ideal cipher. Cryptography and Communications, 7 (4):509–533, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-015-0128-7.

Rahimipour:2015:EML

[120] A. R. Rahimipour, A. R. Ashrafi, and A. Gholami. The existence of minimal logarithmic signatures for the sporadic Suzuki and simple Suzuki groups. Cryptography and Communications, 7 (4):535–542, December 2015. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-015-0129-6.

Wang:2016:FRN

[121] Mingxing Wang, Yupeng Jiang, and Dongdai Lin. Further results on the nonlinearity of maximum-length NFSR feedbacks. Cryptography and Communications, 8(1):1–6, January 2016. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s12095-015-0133-
                                        x.

Alahmadi:2016:LMC

[122] Adel Alahmadi, Hussain Alhazmi, Tor Helleseth, Rola Hijazi, Najat Muthana, and Patrick Solé. On the lifted Melas code. Cryptography and Communications, 8(1):7–18, January 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0135-8.

Paterson:2016:ESD

[123] Maura B. Paterson, Douglas R. Stinson, and Yongge Wang. On encoding symbol degrees of array BP-XOR codes. Cryptography and Communications, 8 (1):19–32, January 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0134-9.

Wang:2016:LCD

[124] Qiuyan Wang, Yupeng Jiang, and Dongdai Lin. Linear complexity of Ding-Helleseth sequences of order 2 over GF(l). Cryptography and Communications, 8(1):33–49, January 2016. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-015-0138-
5.

Gao:2016:CRT

[125] Jian Gao, Linzhi Shen, and Fang-Wei Fu. A Chinese remainder theorem approach to skew generalized quasi-cyclic codes over finite fields. Cryptography and Communications, 8 (1):51–66, January 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0140-y.

Cusick:2016:CEC

[126] Thomas W. Cusick and Pantelimon Stănică. Counting equivalence classes for monomial rotation symmetric Boolean functions with prime dimension. Cryptography and Communications, 8(1):67–81, January 2016. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-015-0143-
8.

Sharma:2016:CSD

[127] Anuradha Sharma and Amit K. Sharma. Construction of self-dual codes over Z_{2^{m}}. Cryptography and Communications, 8(1):83–101, January 2016. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-015-0139-
4.

Lv:2016:NDP

[128] Chuan Lv, Tongjiang Yan, and Guozhen Xiao. New developments in q-polynomial codes. Cryptography and Communications, 8(1):103– 112, January 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0147-4.

Banik:2016:CDC

[129] Subhadeep Banik. Conditional differential cryptanalysis of 105 round Grain v1. Cryptography and Communications, 8(1):113–137, January 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0146-5.

Honkala:2016:EOI

[130] Iiro Honkala, Olivier Hudry, and Antoine Lobstein. On the ensemble of optimal identifying codes in a twin-free graph. Cryptography and Communications, 8(1):139–153, January 2016. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-015-0148-
3.

Xu:2016:CNA

[131] Guangkui Xu, Xiwang Cao, and Shanding Xu. Constructing new APN functions and bent functions over finite fields of odd characteristic via the switching method. Cryptography and Communications, 8(1):155– 171, January 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0145-6.

Budaghyan:2016:ESI

[132] Lilya Budaghyan, Tor Helleseth, and Alexander Kholosha. Editorial: Special issue on Boolean functions and their applications. Cryptography and Communications, 8(2):173–174, April 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s12095-015-0171-4.pdf.

Katz:2016:NOP

[133] Daniel J. Katz and Philippe Langevin.

New open problems related to old conjectures by Helleseth. Cryptography and Communications, 8(2):175– 189, April 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0137-6.

Kasikci:2016:SCQ

[134] Canan Kaşıkcı, Wilfried Meidl, and Alev Topuzoğlu. Spectra of a class of quadratic functions: Average behaviour and counting functions. Cryptography and Communications, 8(2): 191–214, April 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-015-0142-9.

Li:2016:QPL

[135] Chunlei Li and Tor Helleseth. Quasiperfect linear codes from planar and APN functions. Cryptography and Communications, 8(2):215–227, April 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s12095-015-0132-
                                        y.

Mesnager:2016:FCI

[136] Sihem Mesnager. Further constructions of infinite families of bent functions from new permutations and their duals. Cryptography and Communications, 8(2):229–246, April 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-015-0144-7.

Ivanov:2016:RGA

[137] Georgi Ivanov, Nikolay Nikolov, and Svetla Nikova. Reversed genetic algorithms for generation of bijective Sboxes with good cryptographic properties. Cryptography and Communications, 8(2):247–276, April 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0170-5.

Kazymyrov:2016:IAM

[138] Oleksandr Kazymyrov, Roman Oliynykov, and Håvard Raddum. Influence of addition modulo 2^{n} on algebraic attacks. Cryptography and Communications, 8(2):277–289, April 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0136-7.

Tan:2016:ECD

[139] Yin Tan, Guang Gong, and Bo Zhu.

Enhanced criteria on differential uniformity and nonlinearity of cryptographically significant functions. Cryptography and Communications, 8(2): 291–311, April 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0141-x.

Boyar:2016:VNM

[140] Joan Boyar, Magnus Gausdal Find, and René Peralta. On various nonlinearity measures for boolean functions. Cryptography and Communications, 8(3):313–330, July 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0150-9.

Bogos:2016:SPU

[141] Sonia Bogos, Florian Tramèr, and Serge Vaudenay. On solving LPN using BKW and variants: implementation and analysis. Cryptography and Communications, 8(3):331–369, July 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0149-2.

Cheng:2016:DPS

[142] Yuan Cheng, Wen-Feng Qi, Qun-Xiong Zheng, and Dong Yang. On the distinctness of primitive sequences over Z/(p^{e}q) modulo 2. Cryptography and Communications, 8(3): 371–381, July 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0151-8.

Gu:2016:SPH

[143] Ting Gu and Andrew Klapper. Statistical properties of half-ℓ-sequences. Cryptography and Communications, 8 (3):383–400, July 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0152-7.

Chakraborty:2016:AGA

[144] Kaushik Chakraborty and Subhamoy Maitra. Application of Grover’s algorithm to check non-resiliency of a Boolean function. Cryptography and Communications, 8(3):401–413, July 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s12095-015-0156-3.

Klove:2016:CSL

[145] Torleiv Kløve. On covering sets for limited-magnitude errors. Cryptography and Communications, 8(3): 415–433, July 2016. CODEN ????

ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0154-5.

Lu:2016:WTC

[146] Yi Lu and Yvo Desmedt. Walsh transforms and cryptographic applications in bias computing. Cryptography and Communications, 8(3): 435–453, July 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0155-4.

Chakraborty:2016:MOB

[147] Debrup Chakraborty and Palash Sarkar.

On modes of operations of a block cipher for authentication and authenticated encryption. Cryptography and Communications, 8(4):455– 511, October 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0153-6.

Li:2016:DEJ

[148] Fulin Li, Shixin Zhu, Honggang Hu, and Ting Jiang. Determining the kerror joint linear complexity spectrum for a binary multisequence with period p^{n}. Cryptography and Communications, 8(4):513–523, October 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0157-2.

Xiang:2016:LCG

[149] Can Xiang. Linear codes from a generic construction. Cryptography and Communications, 8(4):525– 539, October 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0158-1.

Xu:2016:OAC

[150] Guangkui Xu, Xiwang Cao, and Shanding Xu. Optimal p-ary cyclic codes with minimum distance four from monomials. Cryptography and Communications, 8(4):541–554, October 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0159-0.

Dravie:2016:MRV

[151] Brandon Dravie, Jérémy Parriaux, Philippe Guillot, and Gilles Millérioux. Matrix representations of vectorial Boolean functions and eigenanalysis. Cryptography and Communications, 8 (4):555–577, October 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0160-7.

Chin:2016:TBS

[152] Ji-Jian Chin, Syh-Yuan Tan, Swee-Huay Heng, and Raphael C.-W. Phan. Twin-Beth: Security under active and concurrent attacks for the Beth identitybased identification scheme. Cryptography and Communications, 8(4): 579–591, October 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0162-5.

Jang:2016:LLP

[153] Ji-Woong Jang and Dae-Woon Lim.

Large low probability of intercept properties of the quaternary sequence with optimal correlation property constructed by Legendre sequences. Cryptography and Communications, 8(4): 593–604, October 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0161-6.

Wang:2016:GCN

[154] Qiuyan Wang and Dongdai Lin. Generalized cyclotomic numbers of order two and their applications. Cryptography and Communications, 8(4): 605–616, October 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0165-2.

Sharma:2016:CCF

[155] Anuradha Sharma and Saroj Rani. On constacyclic codes over finite fields. Cryptography and Communications, 8 (4):617–636, October 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0163-4.

Schmidt:2016:NMR

[156] Kai-Uwe Schmidt. Nonlinearity measures of random Boolean functions. Cryptography and Communications, 8 (4):637–645, October 2016. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0164-3.

Ding:2017:PSF

[157] Cunsheng Ding and Zhengchun Zhou. Preface: Special functions and codes. Cryptography and Communications, 9 (1):1–2, January 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s12095-016-0201-x.pdf.

Abdukhalikov:2017:BFL

[158] Kanat Abdukhalikov and Sihem Mesnager. Bent functions linear on elements of some classical spreads and presemifields spreads. Cryptography and Communications, 9(1):3– 21, January 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-016-0195-4.

Bezzateev:2017:QCG

[159] Sergey Bezzateev and Natalia Shekhunova. Quasi-cyclic Goppa codes with special Goppa polynomials and matched location sets. Cryptography and Communications, 9(1):23–39, January 2017. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s12095-016-0196-
                                        3.

Heng:2017:OCA

[160] Ziling Heng and Qin Yue. Optimal codebooks achieving the Levenshtein bound from generalized bent functions over Z_{4}. Cryptography and Communications, 9(1):41–53, January 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-016-0194-5.

Li:2017:CCL

[161] Chengju Li, Sunghan Bae, and Haode Yan. A construction of codes with linearity from two linear codes. Cryptography and Communications, 9(1): 55–69, January 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-016-0193-6.

Mesnager:2017:LCF

[162] Sihem Mesnager. Linear codes with few weights from weakly regular bent functions based on a generic construction. Cryptography and Communications, 9 (1):71–84, January 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-016-0186-5.

Wang:2017:SCC

[163] Qi Wang. Some cyclic codes with prime length from cyclotomy of order 4. Cryptography and Communications, 9(1):85–92, January 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-016-0188-3.

Xiang:2017:CLC

[164] Can Xiang, Chunming Tang, and Keqin Feng. A class of linear codes with a few weights. Cryptography and Communications, 9(1): 93–116, January 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-016-0200-y.

Xu:2017:TCA

[165] Guangkui Xu, Xiwang Cao, and Shanding Xu. Two classes of pary bent functions and linear codes with three or four weights. Cryptography and Communications, 9(1): 117–131, January 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-016-0199-0.

Yang:2017:CTW

[166] Shudi Yang, Zheng-An Yao, and Chang-An Zhao. A class of threeweight linear codes and their complete weight enumerators. Cryptography and Communications, 9(1): 133–149, January 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-016-0187-4.

Zhang:2017:CWE

[167] Dan Zhang, Cuiling Fan, Daiyuan Peng, and Xiaohu Tang. Complete weight enumerators of some linear codes from quadratic forms. Cryptography and Communications, 9(1): 151–163, January 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-016-0190-9.

Zhou:2017:CNF

[168] Yue Zhou and Longjiang Qu. Constructions of negabent functions over finite fields. Cryptography and Communications, 9(2):165–180, March 2017. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-015-0167-
0.

Li:2017:SRG

[169] Xin Li and Zihui Liu. On the second relative greedy weight. Cryptography and Communications, 9(2): 181–197, March 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0168-z.

Wang:2017:CDG

[170] Lisha Wang and Xiaohu Tang. On the correlation distribution of the generalized maximal length Z_{4}-sequences. Cryptography and Communications, 9 (2):199–215, March 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0169-y.

Salagean:2017:CCF

[171] Ana Sălăgean and Matei Mandache-Sălăgean. Counting and characterising functions with “fast points” for differential attacks. Cryptography and Communications, 9(2):217–239, March 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s12095-015-0166-1.pdf.

Bandi:2017:MFN

[172] Rama Krishna Bandi, Maheshanand Bhaintwal, and Nuh Aydin. A mass formula for negacyclic codes of length 2^{k} and some good negacyclic codes over Z_{4} + uZ_{4}. Cryptography and Communications, 9(2): 241–272, March 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0172-3. See cor-

rections [430].

Dubrova:2017:ESC

[173] Elena Dubrova and Martin Hell. Espresso: A stream cipher for 5G wireless communication systems. Cryptography and Communications, 9(2): 273–289, March 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0173-2.

Wu:2017:GGQ

[174] Tingting Wu, Jian Gao, and Fang-Wei Fu. 1-generator generalized quasi-cyclic codes over Z_{4}. Cryptography and Communications, 9(2): 291–299, March 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0175-0.

Gangopadhyay:2017:CBF

[175] Sugata Gangopadhyay, Aditi Kar Gangopadhyay, Spyridon Pollatos, and Pantelimon Stănică. Cryptographic Boolean functions with biased inputs. Cryptography and Communications, 9 (2):301–314, March 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0174-1.

Wang:2017:KTW

[176] Qiuyan Wang, Kelan Ding, Dongdai Lin, and Rui Xue. A kind of three-weight linear codes. Cryptography and Communications, 9(3): 315–322, May 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0180-3.

Heng:2017:CWD

[177] Ziling Heng and Qin Yue. Complete weight distributions of two classes of cyclic codes. Cryptography and Communications, 9(3):323–343, May 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s12095-015-0177-
y.

Liu:2017:NBL

[178] Jian Liu, Sihem Mesnager, and Lusheng Chen. On the nonlinearity of Sboxes and linear codes. Cryptography and Communications, 9(3):345– 361, May 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-015-0176-z.

Peng:2017:NDU

[179] Jie Peng and Chik How Tan. New differentially 4-uniform permutations by modifying the inverse function on subfields. Cryptography and Communications, 9(3):363–378, May 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s12095-016-0181-x.

Esmaeili:2017:NCF

[180] M. Esmaeili, M. Moosavi, and T. A. Gulliver. A new class of Fibonacci sequence based error correcting codes. Cryptography and Communications, 9 (3):379–396, May 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-015-0178-x.

Cheng:2017:AOS

[181] Minquan Cheng, Jing Jiang, and Xiaohu Tang. Asymptotically optimal \overline{2}-separable codes with length 4. Cryptography and Communications, 9 (3):397–405, May 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-016-0182-9.

Liang:2017:SNC

[182] Miao Liang, Lijun Ji, and Jingcai Zhang. Some new classes of 2-fold optimal or perfect splitting authentication codes. Cryptography and Communications, 9(3):407–430, May 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s12095-015-0179-9.

Lin:2017:PCR

[183] Zhiqiang Lin, Dongdai Lin, and Dingyi Pei. Practical construction of ring LF-SRs and ring FCSRs with low diffusion delay for hardware cryptographic applications. Cryptography and Communications, 9(4):431–443, July 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Chen:2017:LCT

[184] Zhixiong Chen. Linear complexity and trace representation of quaternary sequences over Z_{4} based on generalized cyclotomic classes modulo pq. Cryptography and Communications, 9(4):445–458, July 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Raka:2017:CCQ

[185] Madhu Raka, Leetika Kathuria, and Mokshi Goyal. (1 − 2u^{3})-constacyclic codes and quadratic residue codes over F_{p}[u]/⟨u^{4} − u⟩. Cryptography and Communications, 9(4):459–473, July 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Kewat:2017:CCS

[186] Pramod Kumar Kewat and Priti Kumari. Cyclic codes from the second class two-prime Whiteman’s generalized cyclotomic sequence with order 6. Cryptography and Communications, 9 (4):475–499, July 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Merai:2017:ECL

[187] László Mérai, Harald Niederreiter, and Arne Winterhof. Expansion complexity and linear complexity of sequences over finite fields. Cryptography and Communications, 9(4):501–509, July 2017. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Han:2017:NSO

[188] Hongyu Han, Daiyuan Peng, and Udaya Parampalli. New sets of optimal lowhit-zone frequency-hopping sequences based on m-sequences. Cryptography and Communications, 9(4):511–522, July 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Sarkar:2017:PSB

[189] Santanu Sarkar, Prakash Dey, Avishek Adhikari, and Subhamoy Maitra. Probabilistic signature based generalized framework for differential fault analysis of stream ciphers. Cryptography and Communications, 9(4):523–543, July 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Wang:2017:CWE

[190] Xianfang Wang, Jian Gao, and Fang-Wei Fu. Complete weight enumerators of two classes of linear codes. Cryptography and Communications, 9(5):545– 562, September 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Xue:2017:BLA

[191] Shuai Xue, Wen-Feng Qi, and Xiao-Yuan Yang. On the best linear approximation of addition modulo 2^{n}. Cryptography and Communications, 9(5):563– 580, September 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Laing:2017:LMS

[192] Thalia M. Laing, Keith M. Martin, Maura B. Paterson, and Douglas R. Stinson. Localised multisecret sharing. Cryptography and Communications, 9 (5):581–597, September 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Cao:2017:CCO

[193] Yuan Cao and Qingguo Li. Cyclic codes of odd length over Z_{4}[u]/⟨u^{k}⟩. Cryptography and Communications, 9(5):599– 624, September 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Wang:2017:SSS

[194] Xianfang Wang, Can Xiang, and Fang-Wei Fu. Secret sharing schemes for compartmented access structures. Cryptography and Communications, 9(5):625– 635, September 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Shi:2017:TTW

[195] Minjia Shi, Rongsheng Wu, Yan Liu, and Patrick Solé. Two and three weight codes over F_{p} + uF_{p}. Cryptography and Communications, 9(5):637– 646, September 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Tang:2017:CHB

[196] Chunming Tang and Yanfeng Qi. A class of hyper-bent functions and Kloosterman sums. Cryptography and Communications, 9(5):647–664, September 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

Alaca:2017:CVS

[197] Saban Alaca and Goldwyn Millar. Character values of the Sidelnikov–Lempel– Cohn–Eastman sequences. Cryptography and Communications, 9(6):665– 682, November 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s12095-016-
0208-3.

Edemskiy:2017:DSH

[198] Vladimir Edemskiy and Xiaoni Du. Design sequences with high linear complexity over finite fields using generalized cyclotomy. Cryptography and Communications, 9(6):683–691, November 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-016-0209-2.

Li:2017:SCP

[199] Nian Li and Tor Helleseth. Several classes of permutation trinomials from Niho exponents. Cryptography and Communications, 9(6):693–705, November 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-016-0210-9.

Zhou:2017:GMC

[200] Limengnan Zhou, Daiyuan Peng, Hongbin Liang, Changyuan Wang, and Hongyu Han. Generalized methods to construct low-hit-zone frequencyhopping sequence sets and optimal constructions. Cryptography and Communications, 9(6):707–728, November 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0211-3.

Wolfmann:2017:SBF

[201] J. Wolfmann. Sequences of bent functions and near-bent functions. Cryptography and Communications, 9(6):729– 736, November 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s12095-017-
0212-2.

Dietrich:2017:NFA

[202] Heiko Dietrich and Nathan Jolly. A new family of arrays with low autocorrelation. Cryptography and Communications, 9(6):737–748, November 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0214-0.

Ferard:2017:IHS

[203] Eric Férard. On the irreducibility of the hyperplane sections of Fermat varieties in P^{3} in characteristic 2. II. Cryptography and Communications, 9(6):749–767, November 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0213-1.

Helleseth:2018:ESI

[204] Tor Helleseth and Bart Preneel. Editorial: Special issue on recent trends in cryptography. Cryptography and Communications, 10(1):1–3, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/article/10.
1007/s12095-017-0269-y;   http://
link.springer.com/content/pdf/10.
1007/s12095-017-0269-y.pdf.

Scott:2018:MTK

[205] Michael Scott. Missing a trick: Karatsuba variations. Cryptography and Communications, 10(1):5–15, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0217-x.

Ferradi:2018:BAM

[206] Houda Ferradi, Rémi Géraud, Diana Maimut, David Naccache, and Hang Zhou. Backtracking-assisted multiplication. Cryptography and Communications, 10(1):17–26, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0254-5.

Bogos:2018:CHE

[207] Sonia Bogos, John Gaspoz, and Serge Vaudenay. Cryptanalysis of a homomorphic encryption scheme. Cryptography and Communications, 10(1): 27–39, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s12095-017-
                                        0243-8.

Lacharite:2018:SBB

[208] Marie-Sarah Lacharité. Security of BLS and BGLS signatures in a multi-user setting. Cryptography and Communications, 10(1):41–58, January 2018. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s12095-017-0253-6;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s12095-017-0253-6.pdf.

Nachef:2018:GAS

[209] Valérie Nachef, Jacques Patarin, and Emmanuel Volte. Generic attacks with standard deviation analysis on a-Feistel schemes. Cryptography and Communications, 10(1):59–77, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0244-7.

Hemenway:2018:ERS

[210] Brett Hemenway and Rafail Ostrovsky.

Efficient robust secret sharing from expander graphs. Cryptography and Communications, 10(1):79–99, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0215-z.

deCherisey:2018:OPM

[211] Éloi de Chérisey, Sylvain Guilley, Annelie Heuser, and Olivier Rioul. On the optimality and practicability of mutual information analysis in some scenarios. Cryptography and Communications, 10(1):101–121, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0241-x.

Rose:2018:KBT

[212] Gregory G. Rose. KISS: A bit too simple. Cryptography and Communications, 10(1):123–137, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/article/10.
1007/s12095-017-0225-x;   http://
link.springer.com/content/pdf/10.
1007/s12095-017-0225-x.pdf.

Poh:2018:SSE

[213] Geong Sen Poh, Moesfa Soeheila Mohamad, and Ji-Jian Chin. Searchable symmetric encryption over multiple servers. Cryptography and Communications, 10(1):139–158, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0232-y.

Fabsic:2018:GIC

[214] Tomás Fabsic, Otokar Grosek, Karol Nemoga, and Pavol Zajac. On generating invertible circulant binary matrices with a prescribed number of ones. Cryptography and Communications, 10(1):159–175, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0239-4.

Forler:2018:PBB

[215] Christian Forler, Eik List, Stefan Lucks, and Jakob Wenzel. POEx: A beyond-birthday-bound-secure online cipher. Cryptography and Communications, 10(1):177–193, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0250-9.

Cui:2018:SIA

[216] Tingting Cui, Huaifeng Chen, Long Wen, and Meiqin Wang. Statistical integral attack on CAST-256 and IDEA. Cryptography and Communications, 10(1):195–209, January 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0245-6.

Pierrot:2018:MBE

[217] Cécile Pierrot and Benjamin Wesolowski. Malleability of the blockchain’s entropy. Cryptography and Communications, 10(1):211–233, January 2018.

CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0264-3.

Anbar:2018:MPF

[218] Nurdagül Anbar and Wilfried Meidl.

Modified planar functions and their components. Cryptography and Communications, 10(2):235–249, March 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0218-9.

Chang:2018:CBS

[219] Zuling Chang, Martianus Frederic Ezerman, San Ling, and Huaxiong Wang. Construction of de Bruijn sequences from product of two irreducible polynomials. Cryptography and Communications, 10(2):251–275, March 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0219-8.

Sok:2018:CCQ

[220] Lin Sok, MinJia Shi, and Patrick Solé. Classification and construction of quaternary self-dual bent functions. Cryptography and Communications, 10(2):277–289, March 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0216-y.

Tan:2018:WDC

[221] Pan Tan, Zhengchun Zhou, Deng Tang, and Tor Helleseth. The weight distribution of a class of two-weight linear codes derived from Kloosterman sums. Cryptography and Communications, 10(2):291–299, March 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0221-1.

Luo:2018:BLC

[222] Gaojun Luo, Xiwang Cao, Shanding Xu, and Jiafu Mi. Binary linear codes with two or three weights from Niho exponents. Cryptography and Communications, 10(2):301–318, March 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0220-2.

Ding:2018:SCC

[223] Cunsheng Ding. A sequence construction of cyclic codes over finite fields. Cryptography and Communications, 10(2):319–341, March 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0222-0.

Goyal:2018:QRC

[224] Mokshi Goyal and Madhu Raka. Quadratic residue codes over the ring F_{p}[u]/⟨u^{m} − u⟩ and their Gray images. Cryptography and Communications, 10(2):343–355, March 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0223-z.

Acevedo:2018:PSQ

[225] Santiago Barrera Acevedo and Heiko Dietrich. Perfect sequences over the quaternions and (4n, 2, 4n, 2n)-relative difference sets in C_{n} × Q_{8}. Cryptography and Communications, 10(2): 357–368, March 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-017-
0224-y.

Klapper:2018:MPS

[226] Andrew Klapper. Matrix parametrized shift registers. Cryptography and Communications, 10(2):369–382, March 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0226-9.

Dubrova:2018:MAB

[227] Elena Dubrova, Mats Näslund, Göran Selander, and Fredrik Lindqvist. Message authentication based on cryptographically secure CRC without polynomial irreducibility test. Cryptography and Communications, 10(2): 383–399, March 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/article/10.
1007/s12095-017-0227-8;   http://
link.springer.com/content/pdf/10.
1007/s12095-017-0227-8.pdf.

Sharma:2018:EFS

[228] Anuradha Sharma and Taranjot Kaur.

Enumeration formulae for self-dual, self-orthogonal and complementarydual quasi-cyclic codes over finite fields. Cryptography and Communications, 10(3):401–435, May 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0228-7.

Xu:2018:TCN

[229] Shanding Xu, Xiwang Cao, Guangkui Xu, and Gaojun Luo. Two classes of near-optimal frequency-hopping sequence sets with prime-power period. Cryptography and Communications, 10(3):437–454, May 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0229-6.

Wang:2018:LBA

[230] Li-Ping Wang and Daqing Wan. On lattice-based algebraic feedback shift registers synthesis for multisequences. Cryptography and Communications, 10(3):455–465, May 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0230-0.

Sun:2018:EAD

[231] Yuhua Sun, Qiang Wang, and Tongjiang Yan. The exact autocorrelation distribution and 2-adic complexity of a class of binary sequences with almost optimal autocorrelation. Cryptography and Communications, 10(3):467–477, May 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0233-x.

Jana:2018:RRK

[232] Amit Jana and Goutam Paul. Revisiting RC4 key collision: Faster search algorithm and new 22-byte colliding key pairs. Cryptography and Communications, 10(3):479–508, May 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0231-z.

Boztas:2018:GNS

[233] Serdar Boztas, Ferruh Özbudak, and Eda Tekin. Generalized nonbinary sequences with perfect autocorrelation, flexible alphabets and new periods. Cryptography and Communications, 10(3):509–517, May 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0237-6.

Sobhani:2018:CCN

[234] R. Sobhani. Cyclic codes over a noncommutative finite chain ring. Cryptography and Communications, 10(3): 519–530, May 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-017-
0238-5.

Li:2018:PPF

[235] Kangquan Li, Longjiang Qu, Xi Chen, and Chao Li. Permutation polynomials of the form cx + Tr_{q^{l}/q}(x^{a}) and permutation trinomials over finite fields with even characteristic. Cryptography and Communications, 10(3):531–554, May 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0236-7.

Zhang:2018:BCO

[236] Xuli Zhang, Jing Jiang, and Minquan Cheng. Bounds and constructions for \overline{3}-strongly separable codes with length 3. Cryptography and Communications, 10(3):555–565, May 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0235-8.

Zha:2018:NRP

[237] Zhengbang Zha, Lei Hu, and Zhizheng Zhang. New results on permutation polynomials of the form (x^{pm} −x+δ)^{s} + x^{pm} + x over Fp^{2m}. Cryptography and Communications, 10(3):567–578, May 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0234-9.

Liu:2018:NQC

[238] Xiusheng Liu, Hai Q. Dinh, Hualu Liu, and Long Yu. On new quantum codes from matrix product codes. Cryptography and Communications, 10(4):579–589, July 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0242-9.

Yi:2018:GMC

[239] Zongxiang Yi, Zhiqiang Lin, and Lishan Ke. A generic method to construct zero-difference balanced functions. Cryptography and Communications, 10(4):591–609, July 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0247-4.

Alsalami:2018:CHA

[240] Yousuf Alsalami. Constructions with high algebraic degree of differentially 4-uniform (n, n − 1)-functions and differentially 8-uniform (n, n − 2)- functions. Cryptography and Communications, 10(4):611–628, July 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0246-5.

Mandal:2018:ITD

[241] Kalikinkar Mandal, Bo Yang, Guang Gong, and Mark Aagaard. On ideal ttuple distribution of filtering de Bruijn sequence generators. Cryptography and Communications, 10(4):629–641, July 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0248-3.

Sun:2018:SPD

[242] Zhonghua Sun, Shixin Zhu, and Liqi Wang. The symbol-pair distance distribution of a class of repeated-root cyclic codes over F_{p^{m}}. Cryptography and Communications, 10(4):643–653, July 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0249-2.

Castro:2018:EDE

[243] Francis N. Castro, Luis A. Medina, and Ivelisse M. Rubio. Exact 2-divisibility of exponential sums associated to boolean functions. Cryptography and Communications, 10(4):655–666, July 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0252-7.

Barcucci:2018:NOC

[244] Elena Barcucci, Antonio Bernini, Stefano Bilotta, and Renzo Pinzani. A 2D non-overlapping code over a q-ary alphabet. Cryptography and Communications, 10(4):667–683, July 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0251-8.

Jitman:2018:GIS

[245] Somphong Jitman. Good integers and some applications in coding theory. Cryptography and Communications, 10(4):685–704, July 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0255-4. See cor-

rection [270].

Tuxanidy:2018:CDS

[246] Aleksandr Tuxanidy and Qiang Wang. Characteristic digit-sum sequences. Cryptography and Communications, 10 (4):705–717, July 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s12095-017-
                                        0256-3.

Galvez:2018:SBB

[247] Lucky Galvez, Jon-Lark Kim, Nari Lee, Young Gun Roe, and Byung-Sun Won. Some bounds on binary LCD codes. Cryptography and Communications, 10(4):719–728, July 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0258-1.

Maitra:2018:GE

[248] Subhamoy Maitra. Guest editorial. Cryptography and Communications, 10 (5):729–730, September 2018. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/article/10.
1007/s12095-018-0312-7;   http://
link.springer.com/content/pdf/10.
1007/s12095-018-0312-7.pdf.

Jha:2018:RBB

[249] Ashwin Jha and Mridul Nandi. On rate-1 and beyond-the-birthday bound secure online ciphers using tweakable block ciphers. Cryptography and Communications, 10(5):731–753, September 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0275-0.

Cui:2018:SID

[250] Tingting Cui, Huaifeng Chen, Sihem Mesnager, Ling Sun, and Meiqin Wang. Statistical integral distinguisher with multi-structure and its application on AES-like ciphers. Cryptography and Communications, 10(5):755– 776, September 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-018-
0286-5.

Paterson:2018:SAC

[251] Kenneth G. Paterson and Jacob C. N. Schuldt. Statistical attacks on cookie masking for RC4. Cryptography and Communications, 10(5): 777–801, September 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/article/10.
1007/s12095-018-0280-y;   http://
link.springer.com/content/pdf/10.
1007/s12095-018-0280-y.pdf.

Hamann:2018:DAS

[252] Matthias Hamann, Matthias Krause, Willi Meier, and Bin Zhang. Design and analysis of small-state grain-like stream ciphers. Cryptography and Communications, 10(5):803–834, September 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0261-6.

Samajder:2018:SPM

[253] Subhabrata Samajder and Palash Sarkar. Success probability of multiple/multidimensional linear cryptanalysis under general key randomisation hypotheses. Cryptography and Communications, 10(5):835–879, September 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0257-2.

Paul:2018:ABP

[254] Goutam Paul and Souvik Ray. Analysis of burn-in period for RC4 state transition. Cryptography and Communications, 10(5):881–908, September 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0287-4.

Carlet:2018:SPS

[255] Claude Carlet and Sylvain Guilley. Statistical properties of side-channel and fault injection attacks using coding theory. Cryptography and Communications, 10(5):909–933, September 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0271-4.

Bhattacharya:2018:NCS

[256] Srimanta Bhattacharya and Mridul Nandi. A note on the chi-square method: A tool for proving cryptographic security. Cryptography and Communications, 10(5):935–957, September 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0276-z.

Hamann:2018:SCP

[257] Matthias Hamann and Matthias Krause.

On stream ciphers with provable beyond-the-birthday-bound security against time-memory-data tradeoff attacks. Cryptography and Communications, 10(5):959–1012, September 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0294-5.

Merai:2018:PAS

[258] László Mérai and Arne Winterhof. On the pseudorandomness of automatic sequences. Cryptography and Communications, 10(6):1013–1022, November 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0260-7.

Bai:2018:NCP

[259] Tao Bai and Yongbo Xia. A new class of permutation trinomials constructed from Niho exponents. Cryptography and Communications, 10(6): 1023–1036, November 2018. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0263-4.

Meidl:2018:NAB

[260] Wilfried Meidl and Ísabel Pirsic. On the normality of p-ary bent functions. Cryptography and Communications, 10 (6):1037–1049, November 2018. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s12095-017-0259-0;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s12095-017-0259-0.pdf.

Carlet:2018:NMB

[261] Claude Carlet. On the nonlinearity of monotone Boolean functions. Cryptography and Communications, 10(6):1051–1061, November 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0262-5.

Gu:2018:CIF

[262] Ting Gu, Zhixiong Chen, and Andrew Klapper. Correlation immune functions with respect to the qtransform. Cryptography and Communications, 10(6):1063–1073, November 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-017-0267-0.

UlHasan:2018:NVP

[263] Sartaj Ul Hasan, Daniel Panario, and Qiang Wang. Nonlinear vectorial primitive recursive sequences. Cryptography and Communications, 10(6):1075– 1090, November 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s12095-017-
                                        0265-2.

Luo:2018:CWE

[264] Gaojun Luo and Xiwang Cao. Complete weight enumerators of three classes of linear codes. Cryptography and Communications, 10(6):1091–1108, November 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0270-5.

Luo:2018:CCM

[265] Rong Luo and Udaya Parampalli. Cyclic codes over M_{2}(F_{2} + uF_{2}). Cryptography and Communications, 10(6):1109– 1117, November 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-017-
0266-1.

Luo:2018:FCO

[266] Gaojun Luo and Xiwang Cao. Five classes of optimal two-weight linear codes. Cryptography and Communications, 10(6):1119–1135, November 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0272-3.

Samajder:2018:MTD

[267] Subhabrata Samajder and Palash Sarkar. Multiple (truncated) differential cryptanalysis: Explicit upper bounds on data complexity. Cryptography and Communications, 10(6): 1137–1163, November 2018. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0268-z.

Shi:2018:SQM

[268] Xueying Shi, Qin Yue, and Yaotsu Chang. Some quantum MDS codes with large minimum distance from generalized Reed–Solomon codes. Cryptography and Communications, 10(6): 1165–1182, November 2018. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0274-1.

Chang:2018:CSL

[269] Zuling Chang, Martianus Frederic Ezerman, San Ling, and Huaxiong Wang. The cycle structure of LFSR with arbitrary characteristic polynomial over finite fields. Cryptography and Communications, 10(6):1183–1202, November 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0273-2.

Jitman:2018:CGI

[270] Somphong Jitman. Correction to: Good integers and some applications in coding theory. Cryptography and Communications, 10(6):1203, November 2018. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/article/10.
1007/s12095-018-0314-5;   http://
link.springer.com/content/pdf/10.
1007/s12095-018-0314-5.pdf.  See

[245].

Budaghyan:2019:ESIa

[271] Lilya Budaghyan, Claude Carlet, and Tor Helleseth. Editorial: Special issue on Boolean functions and their applications. Cryptography and Communications, 11(1):1–2, ???? 2019. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/article/10.
1007/s12095-018-0341-2;   http://
link.springer.com/content/pdf/10.
1007/s12095-018-0341-2.pdf.

Villa:2019:AFL

[272] Irene Villa. On APN functions L_{1}(x^{3}) + L_{2}(x^{9}) with linear L_{1} and L_{2}. Cryptography and Communications, 11(1):3–20, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0283-8.

Idrisova:2019:AGA

[273] Valeriya Idrisova. On an algorithm generating 2-to-1 APN functions and its applications to “the big APN problem”. Cryptography and Communications, 11(1):21–39, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0310-9.

Mariot:2019:CAB

[274] Luca Mariot, Stjepan Picek, Alberto Leporati, and Domagoj Jakobovic. Cellular automata based S-boxes. Cryptography and Communications, 11(1): 41–62, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-018-
0311-8.

Mesnager:2019:NBF

[275] Sihem Mesnager, Zhengchun Zhou, and Cunsheng Ding. On the nonlinearity of Boolean functions with restricted input. Cryptography and Communications, 11(1):63–76, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0293-6.

Mandal:2019:NCA

[276] Bimal Mandal, Pantelimon Stanica, and Sugata Gangopadhyay. New classes of p-ary bent functions. Cryptography and Communications, 11(1):77–92, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0290-9.

Calik:2019:MCV

[277] Çagdas Çalik, Meltem Sönmez Turan, and René Peralta. The multiplicative complexity of 6-variable Boolean functions. Cryptography and Communications, 11(1):93–107, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0297-2.

Boyar:2019:SLD

[278] Joan Boyar, Magnus Gausdal Find, and René Peralta. Small low-depth circuits for cryptographic applications. Cryptography and Communications, 11(1):109–127, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0296-3.

Li:2019:NPT

[279] Nian Li and Tor Helleseth. New permutation trinomials from Niho exponents over finite fields with even characteristic. Cryptography and Communications, 11(1):129–136, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0321-6.

Liu:2019:SNC

[280] Hongwei Liu and Youcef Maouche.

Several new classes of linear codes with few weights. Cryptography and Communications, 11(2):137–146, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0277-y.

Martinez-Penas:2019:USR

[281] Umberto Martínez-Peñas. Universal secure rank-metric coding schemes with optimal communication overheads. Cryptography and Communications, 11(2):147–166, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0279-4.

Wang:2019:IDE

[282] Lin Wang and Zhi Hu. Injectivity on distribution of elements in the compressed sequences derived from primitive sequences over Zpe. Cryptography and Communications, 11(2):167–189, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-017-0278-x.

Michel:2019:SNB

[283] Jerod Michel and Qi Wang. Some new balanced and almost balanced quaternary sequences with low autocorrelation. Cryptography and Communications, 11(2):191–206, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0281-x.

Pasalic:2019:BFN

[284] Enes Pasalic, Samir Hodžić, Fengrong Zhang, and Yongzhuang Wei. Bent functions from nonlinear permutations and conversely. Cryptography and Communications, 11(2):207–225, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0282-9.

Deng:2019:MCP

[285] Huali Deng and Dabin Zheng. More classes of permutation trinomials with Niho exponents. Cryptography and Communications, 11(2):227–236, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0284-7.

Geil:2019:AVC

[286] Olav Geil and Ferruh Özbudak. On affine variety codes from the Klein quartic. Cryptography and Communications, 11(2):237–257, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0285-6.

Yan:2019:NCM

[287] Haode Yan. A note on the constructions of MDS self-dual codes. Cryptography and Communications, 11(2): 259–268, ???? 2019. CODEN ????

ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-018-
0288-3.

Wang:2019:NBC

[288] Qichun Wang and Pantelimon Stanica. New bounds on the covering radius of the second order Reed–Muller code of length 128. Cryptography and Communications, 11(2):269–277, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0289-2.

Li:2019:CIP

[289] Kangquan Li, Longjiang Qu, and Qiang Wang. Compositional inverses of permutation polynomials of the form x^{r}h(x^{s}) over finite fields. Cryptography and Communications, 11(2):279–298, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0292-7.

Gao:2019:UBC

[290] Zhe Gao, Chao Li, and Yue Zhou. Upper bounds and constructions of complete asynchronous channel hopping systems. Cryptography and Communications, 11(2):299–312, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0295-4.

Wu:2019:SCP

[291] Gaofei Wu and Nian Li. Several classes of permutation trinomials over F_{5^{n}} from Niho exponents. Cryptography and Communications, 11(2):313–324, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0291-8.

Ashraf:2019:QCC

[292] Mohammad Ashraf and Ghulam Mohammad. Quantum codes over F_{p} from cyclic codes over F_{p}[u, v]/⟨u^{2} − 1, v^{3} − v, uv − vu⟩. Cryptography and Communications, 11(2):325–335, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0299-0.

Sun:2019:LBA

[293] Yuhua Sun, Qiang Wang, and Tongjiang Yan. A lower bound on the 2-adic complexity of the modified Jacobi sequence. Cryptography and Communications, 11(2):337–349, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0300-y.

Rostami:2019:CWW

[294] Saeed Rostami, Elham Shakour, Mohammad Ali Orumiehchiha, and Josef Pieprzyk. Cryptanalysis of WG-8 and WG-16 stream ciphers. Cryptography and Communications, 11(2): 351–362, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s12095-018-
                                        0298-1.

Budaghyan:2019:ESIb

[295] Lilya Budaghyan, Chunlei Li, and Matthew G. Parker. Editorial: Special issue on mathematical methods for cryptography. Cryptography and Communications, 11(3):363–365, ???? 2019.

CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/article/10.
1007/s12095-019-00356-8;   http:
//link.springer.com/content/pdf/
10.1007/s12095-019-00356-8.pdf.

Nyberg:2019:ALC

[296] Kaisa Nyberg. Affine linear cryptanalysis. Cryptography and Communications, 11(3):367–377, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0325-2.

Nikova:2019:DPF

[297] Svetla Nikova, Ventzislav Nikov, and Vincent Rijmen. Decomposition of permutations in a finite field. Cryptography and Communications, 11(3): 379–384, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-018-
0317-2.

Varici:2019:CBU

[298] Kerem Varici, Svetla Nikova, Ventzislav Nikov, and Vincent Rijmen. Constructions of S-boxes with uniform sharing. Cryptography and Communications, 11(3):385–398, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0345-y.

Csirmaz:2019:SSL

[299] László Csirmaz and Péter Ligeti. Secret sharing on large girth graphs. Cryptography and Communications, 11(3): 399–410, ???? 2019. CODEN ????

ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s12095-018-
                                        0338-x.

Chen:2019:EAW

[300] Yao Chen, Benjamin M. Case, Shuhong Gao, and Guang Gong. Error analysis of weak Poly–LWE instances. Cryptography and Communications, 11(3): 411–426, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s12095-018-
                                        0301-x.

Felke:2019:SBP

[301] Patrick Felke. On the security of biquadratic C^{∗} public-key cryptosystems and its generalizations. Cryptography and Communications, 11(3): 427–442, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s12095-018-
                                        0337-y.

Raddum:2019:FUB

[302] Håvard Raddum and Srimathi Varadharajan. Factorization using binary decision diagrams. Cryptography and Communications, 11(3):443–460, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0304-7.

Zinoviev:2019:CKS

[303] V. A. Zinoviev. On classical Kloosterman sums. Cryptography and Communications, 11(3):461–496, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00357-7.

Bojilov:2019:CNF

[304] Assen Bojilov, Lyubomir Borissov, and Yuri Borissov. Computing the number of finite field elements with prescribed absolute trace and cotrace. Cryptography and Communications, 11(3):497–507, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0336-z.

Li:2019:SAN

[305] Nian Li and Xiangyong Zeng. A survey on the applications of Niho exponents. Cryptography and Communications, 11(3):509–548, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0305-6.

Klove:2019:CLT

[306] Torleiv Kløve. Codes of length two correcting single errors of limited size. Cryptography and Communications, 11(3):549–555, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0327-0.

Buratti:2019:HPD

[307] Marco Buratti. Hadamard partitioned difference families and their descendants. Cryptography and Communications, 11(4):557–562, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0308-3.

Tu:2019:CPT

[308] Ziran Tu and Xiangyong Zeng. A class of permutation trinomials over finite fields of odd characteristic. Cryptography and Communications, 11(4):563–583, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0307-4.

Napp:2019:PSS

[309] Diego Napp, Ricardo Pereira, Raquel Pinto, and Paula Rocha. Periodic statespace representations of periodic convolutional codes. Cryptography and Communications, 11(4):585–595, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0313-6.

Harada:2019:SES

[310] Masaaki Harada. Singly even self-dual codes of length 24k + 10 and minimum weight 4k + 2. Cryptography and Communications, 11(4):597–608, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0303-8.

Yang:2019:CWE

[311] Shudi Yang, Qin Yue, Yansheng Wu, and Xiangli Kong. Complete weight enumerators of a class of two-weight linear codes. Cryptography and Communications, 11(4):609–620, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0315-4.

Junnila:2019:OBC

[312] Ville Junnila, Tero Laihonen, and Gabrielle Paris. Optimal bounds on codes for location in circulant graphs. Cryptography and Communications, 11(4):621–640, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0316-3.

Kavut:2019:RSB

[313] Selçuk Kavut and Sevdenur Baloglu. Results on symmetric S-boxes constructed by concatenation of RSSBs. Cryptography and Communications, 11(4): 641–660, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-018-
0318-1.

Zhou:2019:FHS

[314] Limengnan Zhou, Daiyuan Peng, Xing Liu, Hongyu Han, and Zheng Ma. Frequency-hopping sequence sets with good aperiodic Hamming correlation property. Cryptography and Communications, 11(4):661–675, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0320-7.

Harada:2019:BLC

[315] Masaaki Harada and Ken Saito.

Binary linear complementary dual codes. Cryptography and Communications, 11(4):677–696, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0319-0.

Dey:2019:SMR

[316] Sabyasachi Dey and Santanu Sarkar. Settling the mystery of Z_{r} = r in RC4. Cryptography and Communications, 11(4):697–715, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0323-4.

Qian:2019:SDL

[317] Liqin Qian, Minjia Shi, and Patrick Solé. On self-dual and LCD quasitwisted codes of index two over a special chain ring. Cryptography and Communications, 11(4):717–734, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0322-5.

Xiao:2019:BSP

[318] Zibi Xiao, Xiangyong Zeng, Chaoyun Li, and Yupeng Jiang. Binary sequences with period N and nonlinear complexity N − 2. Cryptography and Communications, 11(4):735–757, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0324-3.

Wang:2019:DCC

[319] Gang Wang, Min-Yao Niu, and Fang-Wei Fu. Deterministic constructions of compressed sensing matrices based on codes. Cryptography and Communications, 11(4):759–775, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0328-z.

Oblaukhov:2019:LBS

[320] Alexey Oblaukhov. A lower bound on the size of the largest metrically regular subset of the Boolean cube. Cryptography and Communications, 11(4):777–791, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0326-1.

Gorodilova:2019:DEA

[321] Anastasiya Gorodilova. On the differential equivalence of APN functions. Cryptography and Communications, 11(4):793–813, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0329-y.

Li:2019:CNO

[322] Yu Li, Tongjiang Yan, and Chuan Lv.

Construction of a near-optimal quasicomplementary sequence set from almost difference set. Cryptography and Communications, 11(4):815–824, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0330-5.

Luo:2019:TCA

[323] Gaojun Luo and Xiwang Cao. Two constructions of asymptotically optimal codebooks. Cryptography and Communications, 11(4):825–838, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0331-4.

Climent:2019:CAN

[324] Joan-Josep Climent, Verónica Requena, and Xaro Soler-Escrivà. A construction of Abelian non-cyclic orbit codes. Cryptography and Communications, 11(5):839–852, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0306-5.

Wu:2019:BFS

[325] Gaofei Wu and Matthew Geoffrey Parker. On Boolean functions with several flat spectra. Cryptography and Communications, 11(5):853–880, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0309-2.

Laaksonen:2019:NLB

[326] Antti Laaksonen and Patric R. J. Östergård. New lower bounds on q-ary error-correcting codes. Cryptography and Communications, 11(5): 881–889, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-018-
0302-9.

Aydin:2019:NLC

[327] Nuh Aydin, Ghada Bakbouk, and Jonathan G. G. Lambrinos. New linear codes over non-prime fields. Cryptography and Communications, 11(5): 891–902, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-018-
0333-2.

Bartoli:2019:NBL

[328] Daniele Bartoli, Alexander A. Davydov, Massimo Giulietti, Stefano Marcugini, and Fernanda Pambianco. New bounds for linear codes of covering radii 2 and 3. Cryptography and Communications, 11(5):903–920, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0335-0.

Uguz:2019:RCT

[329] Muhiddin Uguz, Ali Doganaksoy, Fatih Sulak, and Onur Koçak. R-2 composition tests: a family of statistical randomness tests for a collection of binary sequences. Cryptography and Communications, 11(5):921–949, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0334-1.

Wang:2019:NEI

[330] Lin Wang and Zhi Hu. New explicit injective compressing mappings on primitive sequences over Z_{p^{e}}. Cryptography and Communications, 11(5):951–963, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0332-3.

Rabii:2019:NCS

[331] Hila Rabii and Osnat Keren. A new class of security oriented error correcting robust codes. Cryptography and Communications, 11(5):965–978, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0340-3.

Gao:2019:SBR

[332] Ying Gao, Zihui Liu, and Yiwei Liu. The separation of binary relative three-weight codes and its applications. Cryptography and Communications, 11(5):979–992, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0339-9.

Hofer:2019:ONC

[333] Richard Hofer and Arne Winterhof. rth order nonlinearity, correlation measure and least significant bit of the discrete logarithm. Cryptography and Communications, 11(5):993–997, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0344-z.

Naghipour:2019:NCQ

[334] Avaz Naghipour. New classes of quantum codes on closed orientable surfaces. Cryptography and Communications, 11(5):999–1008, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0347-9.

Johnsen:2019:RPE

[335] Trygve Johnsen and Hugues Verdure. Relative profiles and extended weight polynomials of almost affine codes. Cryptography and Communications, 11(5):1009–1020, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0348-8.

Garcia:2019:NMD

[336] Ismael Gutiérrez García, Daladier Jabba Molinares, and Ivan Molina Naizir. A novel maximum distance separable code to generate universal identifiers. Cryptography and Communications, 11(5):1021–1035, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0346-x.

Liang:2019:LCD

[337] Yana Liang, Jiali Cao, Xingfa Chen, Shiping Cai, and Xiang Fan. Linear complexity of Ding–Helleseth generalized cyclotomic sequences of order eight. Cryptography and Communications, 11(5):1037–1056, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-018-0343-0.

Wang:2019:FRP

[338] Libo Wang, Baofeng Wu, Xiaoqiang Yue, and Yanbin Zheng. Further results on permutation trinomials with Niho exponents. Cryptography and Communications, 11(5):1057–1068, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-0349-2.

Radonjic:2019:ICC

[339] Aleksandar Radonjic and Vladimir Vujicic. Integer codes correcting sparse byte errors. Cryptography and Communications, 11(5):1069–1077, ???? 2019.

CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-0350-9.

Hodzic:2019:IML

[340] S. Hodžić, E. Pasalic, and A. Chattopadhyay. An iterative method for linear decomposition of index generating functions. Cryptography and Communications, 11(5):1079–1102, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-0351-8.

Guneri:2019:CCL

[341] Cem Güneri, Ferruh Özbudak, and Elif Saçikara. A concatenated construction of linear complementary pair of codes. Cryptography and Communications, 11(5):1103–1114, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-0354-5.

Beemer:2019:ASC

[342] Allison Beemer, Kathryn Haymaker, and Christine A. Kelley. Absorbing sets of codes from finite geometries. Cryptography and Communications, 11(5):1115–1131, ???? 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-0353-6.

Tang:2019:RAB

[343] Chunming Tang, Yanfeng Qi, and Dongmei Huang. Regular p-ary bent functions with five terms and Kloosterman sums. Cryptography and Communications, 11(5):1133–1144, ???? 2019.

CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-0355-4.

Budaghyan:2019:ESI

[344] Lilya Budaghyan, Claude Carlet, and Tor Helleseth. Editorial: Special issue on Boolean functions and their applications 2018. Cryptography and Communications, 11(6):1145–1146, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:

//link.springer.com/article/10.
1007/s12095-019-00409-y;   https:
//link.springer.com/content/pdf/
10.1007/s12095-019-00409-y.pdf.

Canteaut:2019:IGB

[345] Anne Canteaut, Léo Perrin, and Shizhu Tian. If a generalised butterfly is APN then it operates on 6 bits. Cryptography and Communications, 11(6):1147–1164, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00361-x.

Kaleyski:2019:CAF

[346] Nikolay S. Kaleyski. Changing APN functions at two points. Cryptography and Communications, 11(6):1165– 1184, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-019-
00366-6.

Tang:2019:FWA

[347] Deng Tang and Jian Liu. A family of weightwise (almost) perfectly balanced boolean functions with optimal algebraic immunity. Cryptography and Communications, 11(6):1185–1197, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-019-00374-6.

Hou:2019:CPT

[348] Xiang dong Hou. On a class of permutation trinomials in characteristic 2. Cryptography and Communications, 11(6):1199–1210, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-018-0342-1.

Gravel:2019:USP

[349] Claude Gravel, Daniel Panario, and David Thomson. Unicyclic strong permutations. Cryptography and Communications, 11(6):1211–1231, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-019-00384-4.

Meidl:2019:GBF

[350] Wilfried Meidl and Alexander Pott. Generalized bent functions into Z_{p^{k}} from the partial spread and the Maiorana–McFarland class. Cryptography and Communications, 11(6):1233– 1245, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s12095-019-
                                        00370-w.

Mesnager:2019:MCT

[351] Sihem Mesnager, Constanza Riera, and Pantelimon Stanica. Multiple characters transforms and generalized Boolean functions. Cryptography and Communications, 11(6):1247–1260, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00383-5.

Luo:2019:SNC

[352] Gaojun Luo, Xiwang Cao, and Sihem Mesnager. Several new classes of selfdual bent functions derived from involutions. Cryptography and Communications, 11(6):1261–1273, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00371-9.

Cepak:2019:FLT

[353] N. Cepak, E. Pasalic, and A. Muratović-Ribić. Frobenius linear translators giving rise to new infinite classes of permutations and bent functions. Cryptography and Communications, 11(6):1275–1295, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00395-1.

Ozbudak:2019:SRG

[354] Ferruh Özbudak and RumI MelIh Pelen. Strongly regular graphs arising from non-weakly regular bent functions. Cryptography and Communications, 11(6):1297–1306, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00394-2.

Cesarz:2019:ISR

[355] Patrick G. Cesarz and Robert S. Coulter. Image sets with regularity of differences. Cryptography and Communications, 11(6):1307–1337, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-019-00358-6.

Brandao:2019:UBM

[356] Luís T. A. N. Brandão, Çagdas Çalik, Meltem Sönmez Turan, and René Peralta. Upper bounds on the multiplicative complexity of symmetric Boolean functions. Cryptography and Communications, 11(6):1339–1362, November 2019. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-019-00377-3.

Gao:2020:ASQ

[357] Yun Gao, Weijun Fang, and Fang-Wei Fu. On the algebraic structure of quasi-cyclic codes of index 1\frac{1}{2}. Cryptography and Communications, 12(1):1–18, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-019-0352-7.

Mesnager:2020:MTW

[358] Sihem Mesnager, Kwang Ho Kim, Junyop Choe, and Chunming Tang. On the Menezes–Teske–Weng conjecture. Cryptography and Communications, 12(1):19–27, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-019-00359-5.

Shi:2020:PII

[359] Zexia Shi and Fang-Wei Fu. The primitive idempotents of irreducible constacyclic codes and LCD cyclic codes. Cryptography and Communications, 12(1):29–52, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00362-w.

Shi:2020:SDL

[360] Minjia Shi, Hongwei Zhu, Liqin Qian, Lin Sok, and Patrick Solé. On selfdual and LCD double circulant and double negacirculant codes over F_{q} + uF_{q}. Cryptography and Communications, 12(1):53–70, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00363-9.

Milshtein:2020:NTE

[361] Moshe Milshtein. A new twoerror-correcting binary code of length 16. Cryptography and Communications, 12(1):71–75, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00365-7.

Tang:2020:NLB

[362] Deng Tang, Haode Yan, Zhengchun Zhou, and Xiaosong Zhang. A new lower bound on the second-order nonlinearity of a class of monomial bent functions. Cryptography and Communications, 12(1):77–83, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00360-y.

Budaghyan:2020:RBC

[363] L. Budaghyan, Marco Calderini, and I. Villa. On relations between CCZand EA-equivalences. Cryptography and Communications, 12(1):85–100, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s12095-019-00367-5.

Alahmadi:2020:CDM

[364] Adel Alahmadi, Cem Güneri, Buket Özkaya, Hatoon Shoaib, and Patrick Solé. On complementary dual multinegacirculant codes. Cryptography and Communications, 12(1):101–113, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s12095-019-
                                        00364-8.

Roy:2020:FSS

[365] Suman Roy and Srinivasan Krishnaswamy. On the frequency of symbols in sequences generated by nonlinear feedforward generators. Cryptography and Communications, 12(1): 115–126, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s12095-019-
                                        00379-1.

Dougherty:2020:QBC

[366] Steven T. Dougherty, Joseph Gildea, and Abidin Kaya. Quadruple bordered constructions of self-dual codes from group rings. Cryptography and Communications, 12(1):127–146, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00380-8.

Li:2020:DDP

[367] Fengwei Li, Qin Yue, and Yansheng Wu. Designed distances and parameters of new LCD BCH codes over finite fields. Cryptography and Communications, 12(1):147–163, January 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00385-3.

Niu:2020:NCI

[368] Tailin Niu, Kangquan Li, Longjiang Qu, and Qiang Wang. New constructions of involutions over finite fields. Cryptography and Communications, 12(2):165–185, March 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00386-2.

Guo:2020:FNB

[369] Guanmin Guo, Ruihu Li, Yang Liu, and Junli Wang. A family of negacyclic BCH codes of length n = \frac{q^{2m}−1}{2}. Cryptography and Communications, 12(2):187–203, March 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00387-1.

Shuai:2020:DUC

[370] Li Shuai, Lina Wang, Li Miao, and Xianwei Zhou. Differential uniformity of the composition of two functions. Cryptography and Communications, 12(2):205–220, March 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00382-6.

Radonjic:2020:ICC

[371] Aleksandar Radonjic and Vladimir Vujicic. Integer codes correcting burst asymmetric within a byte and double asymmetric errors. Cryptography and Communications, 12(2):221–230, March 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00388-0.  See

correction [372].

Radonjic:2020:CIC

[372] Aleksandar Radonjic and Vladimir Vujicic. Correction to: Integer codes correcting burst asymmetric errors within a byte and double asymmetric errors. Cryptography and Communications, 12 (2):231, March 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-019-
00398-y;   http://link.springer.
com/content/pdf/10.1007/s12095-
019-00398-y.pdf. See [371].

Wang:2020:CSN

[373] Zhongxiao Wang, Qunxiong Zheng, and Wenfeng Qi. The cycle structure of NFSR(f^{d}) and its applications. Cryptography and Communications, 12(2):233–252, March 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00392-4.

Yao:2020:MZZ

[374] Ting Yao and Shixin Zhu. Z_{p}Z_{p^{s}}- additive cyclic codes are asymptotically good. Cryptography and Communications, 12(2):253–264, March 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00397-z.

Sun:2020:CCB

[375] Zhonghua Sun, Shixin Zhu, and Liqi Wang. A class of constacyclic BCH codes. Cryptography and Communications, 12(2):265–284, March 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00401-6.

Araya:2020:MWB

[376] Makoto Araya and Masaaki Harada.

On the minimum weights of binary linear complementary dual codes. Cryptography and Communications, 12(2): 285–300, March 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s12095-019-
00402-5.

Cao:2020:CCS

[377] Yuan Cao and Yonglin Cao. Complete classification for simple root cyclic codes over the local ring Z_{4}[v]/⟨v^{2} + 2v⟩. Cryptography and Communications, 12(2):301–319, March 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL

http://link.springer.com/article/
10.1007/s12095-019-00403-4.

Helleseth:2020:ESI

[378] Tor Helleseth, Wai Ho Mow, and Zhengchun Zhou. Editorial: Special issue on sequences and their applications 2018. Cryptography and Communications, 12(3):321–323, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00433-
                                        3.

Zhang:2020:PZC

[379] Dan Zhang, Matthew Geoffrey Parker, and Tor Helleseth. Polyphase zero correlation zone sequences from generalised bent functions. Cryptography and Communications, 12(3):325–335, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-019-00413-
                                        2.

Krengel:2020:OCP

[380] Evgeny Krengel. One construction of perfect ternary sequences. Cryptography and Communications, 12(3):337–347, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-019-00414-
                                        1.

Zhou:2020:TCQ

[381] Yajing Zhou, Zhengchun Zhou, and Yong Wang. Two constructions for 16-QAM complementary sequence sets with non-power-of-two length. Cryptography and Communications, 12(3):349–362, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-019-00406-
                                        1.

Yang:2020:NQS

[382] Yang Yang and Chunlei Li. New quaternary sequences with optimal oddperiodic autocorrelation magnitude. Cryptography and Communications, 12 (3):363–374, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-019-
00407-0.

Tang:2020:NMB

[383] Deng Tang and Xia Li. A note on the minimal binary linear code. Cryptography and Communications, 12(3): 375–388, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-019-
00412-3.

Gomez-Perez:2020:LFS

[384] Domingo Gómez-Pérez, Ana I. Gómez, and Andrew Tirkel. Large families of sequences for CDMA, frequency hopping, and UWB. Cryptography and Communications, 12(3):389–403, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00399-
x.

Han:2020:DSF

[385] Hongyu Han, Sheng Zhang, and Xing Liu. Decimated m-sequences families with optimal partial Hamming correlation. Cryptography and Communications, 12(3):405–413, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00400-

Sun:2020:TMO

[386] Zhimin Sun, Xiangyong Zeng, and Da Lin. On the N-th maximum order complexity and the expansion complexity of a Rudin–Shapirolike sequence. Cryptography and Communications, 12(3):415–426, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-019-00396-
                                        0.

Li:2020:TCB

[387] Chunlei Li and Yang Yang. On three conjectures of binary sequences with low odd-periodic autocorrelation. Cryptography and Communications, 12(3): 427–442, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-019-
                                        00393-3.

Shi:2020:OWT

[388] Minjia Shi, Chenchen Wang, and Yaoqiang Chang. One-weight and twoweight Z_{2}Z_{2}[u, v] -additive codes. Cryptography and Communications, 12(3): 443–454, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-019-
                                        00391-5.

Mesnager:2020:GHB

[389] Sihem Mesnager. On generalized hyperbent functions. Cryptography and Communications, 12(3):455–468, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s12095-019-00390-
6.

Li:2020:DCP

[390] Shuxing Li and Alexander Pott. A direct construction of primitive formally dual pairs having subsets with unequal sizes. Cryptography and Communications, 12(3):469–483, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00389-
z.

Zhou:2020:FHS

[391] Limengnan Zhou, Hongyu Han, and Xing Liu. Frequency-hopping sequence sets with no-hit-zone through Cartesian product. Cryptography and Communications, 12(3):485–497, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00381-
7.

Egan:2020:PTF

[392] Malcolm Egan. Properties of tight frames that are regular schemes. Cryptography and Communications, 12(3): 499–510, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-019-
00378-2.

Wu:2020:BTS

[393] Yansheng Wu, Qin Yue, and Xiaomeng Zhu. Binary and ternary sequences with a few cross correlations. Cryptography and Communications, 12(3):511–525, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-019-00376-
                                        4.

Budaghyan:2020:PAB

[394] Lilya Budaghyan, Nikolay S. Kaleyski, and Pantelimon Stanica. Partially APN Boolean functions and classes of functions that are not APN infinitely often. Cryptography and Communications, 12(3):527–545, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-019-00372-
                                        8.

Cao:2020:TDC

[395] Xiwang Cao, Gaojun Luo, and Guangkui Xu. Three deterministic constructions of compressed sensing matrices with low coherence. Cryptography and Communications, 12(3):547–558, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-019-00375-
                                        5.

Li:2020:FDO

[396] Xia Li, Cuiling Fan, and Xiaoni Du. A family of distance-optimal minimal linear codes with flexible parameters. Cryptography and Communications, 12 (3):559–567, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-019-
                                        00373-7.

Li:2020:CES

[397] Chengju Li, Qin Yue, and Wei Peng. A class of exponential sums and sequence families. Cryptography and Communications, 12(3):569–584, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00368-
4.

Wang:2020:FSC

[398] Zilong Wang, Jinjin Chai, and Guang Gong. The Fourier spectral characterization for the correlation-immune functions over F_{p}. Cryptography and Communications, 12(3):585–595, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00369-
3.

Srikanth:2020:CSA

[399] Ch. Srikanth. Certain sequence of arithmetic progressions and a new key sharing method. Cryptography and Communications, 12(3):597–612, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00416-
z.

Benerjee:2020:NUF

[400] Krishna Gopal Benerjee and Manish K. Gupta. On non-uniform flower codes. Cryptography and Communications, 12(3):613–643, May 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00430-
6.

Liu:2020:NGH

[401] Zihui Liu and Jinliang Wang. Notes on generalized Hamming weights of some classes of binary codes. Cryptography and Communications, 12(4):645–657, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00404-
3.

Mesnager:2020:NRZ

[402] Sihem Mesnager, Kwang Ho Kim, and Myong Song Jo. On the number of the rational zeros of linearized polynomials and the second-order nonlinearity of cubic Boolean functions. Cryptography and Communications, 12(4):659–674, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00410-
5.

Sun:2020:ACC

[403] Yuhua Sun, Tongjiang Yan, and Lianhai Wang. The 2-adic complexity of a class of binary sequences with optimal autocorrelation magnitude. Cryptography and Communications, 12(4):675–683, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00411-
4.

Beierle:2020:DZS

[404] Christof Beierle, Alex Biryukov, and Aleksei Udovenko. On degree-d zerosum sets of full rank. Cryptography and Communications, 12(4):685–710, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00415-
0.

Singh:2020:SSQ

[405] Jasvinder Singh, Manish Gupta, and Jaskarn Singh Bhullar. On the search of smallest QC-LDPC code with girth six and eight. Cryptography and Communications, 12(4):711–723, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00405-
2.

Zhao:2020:ALC

[406] Lu Zhao. About the linear complexity of quaternary sequences with even length. Cryptography and Communications, 12(4):725–741, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00419-
w.

Pang:2020:SNB

[407] Binbin Pang, Shixin Zhu, and Xiaoshan Kai. Some new bounds on LCD codes over finite fields. Cryptography and Communications, 12(4):743–755, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00417-
y.

Armario:2020:ASD

[408] J. A. Armario and D. L. Flannery.

Almost supplementary difference sets and quaternary sequences with optimal autocorrelation. Cryptography and Communications, 12(4):757–768, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00418-
x.

Gildea:2020:DBC

[409] Joe Gildea, Rhian Taylor, and A. Tylyshchak. Double bordered constructions of selfdual codes from group rings over Frobenius rings. Cryptography and Communications, 12(4):769–784, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00420-
3.

Ortiz-Ubarri:2020:NAO

[410] José Ortiz-Ubarri. New asymptotically optimal three-dimensional wavelength/space/time optical orthogonal codes for OCDMA systems. Cryptography and Communications, 12(4): 785–794, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-019-
00422-1.

Hu:2020:MDN

[411] Liqin Hu and Keqin Feng. The minimum distance of new generalisations of the punctured binary Reed– Muller codes. Cryptography and Communications, 12(4):795–808, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-019-00421-
2.

Mesnager:2020:S

[412] Sihem Mesnager, Kwang Ho Kim, and Dae Song Go. Solving x + x^{2^{l}} + ⋯ + x^{2^{ml}} = a over F_{2^{n}}. Cryptography and Communications, 12(4):809–817, July 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00425-
3.

Budaghyan:2020:ESI

[413] Lilya Budaghyan and Tor Helleseth. Editorial: Special issue on Boolean functions and their applications. Cryptography and Communications, 12(5):819–820, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00454-
y.

Calderini:2020:ECK

[414] Marco Calderini. On the EA-classes of known APN functions in small dimensions. Cryptography and Communications, 12(5):821–840, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00427-
1.

Felke:2020:MMS

[415] Patrick Felke. The multivariate method strikes again: New power functions with low differential uniformity in odd characteristic. Cryptography and Communications, 12(5):841–857, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00437-
z.

Musukwa:2020:LSB

[416] A. Musukwa and M. Sala. On the linear structures of balanced functions and quadratic APN functions. Cryptography and Communications, 12(5):859–880, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00431-
                                        5.

Kutsenko:2020:GAS

[417] Aleksandr Kutsenko. The group of automorphisms of the set of self-dual bent functions. Cryptography and Communications, 12(5):881–898, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00438-
                                        y.

Cesmelioglu:2020:VBF

[418] Ayça Çesmelioglu, Wilfried Meidl, and Alexander Pott. Vectorial bent functions in odd characteristic and their components. Cryptography and Communications, 12(5):899–912, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00444-
                                        0.

Kalayci:2020:PPF

[419] Tekgül Kalayci, Henning Stichtenoth, and Alev Topuzoglu. Permutation polynomials and factorization. Cryptography and Communications, 12(5):913– 934, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.

com/article/10.1007/s12095-020-
00446-y.

Calik:2020:BFM

[420] Çagdas Çalik, Meltem Sönmez Turan, and René Peralta. Boolean functions with multiplicative complexity 3 and 4. Cryptography and Communications, 12(5):935–946, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00445-
z.

Zajac:2020:CPS

[421] Pavol Zajac and Matús Jókay. Cryptographic properties of small bijective S-boxes with respect to modular addition. Cryptography and Communications, 12(5):947–963, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00447-
x.

Li:2020:RRP

[422] Nian Li and Sihem Mesnager. Recent results and problems on constructions of linear codes from cryptographic functions. Cryptography and Communications, 12(5):965–986, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00435-
1.

Kadir:2020:DAG

[423] Wrya K. Kadir and Chunlei Li.

On decoding additive generalized twisted Gabidulin codes. Cryptography and Communications, 12(5):987– 1009, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-020-
                                        00449-9.

Ding:2020:CDS

[424] Cunsheng Ding and Chunming Tang. Combinatorial t-designs from special functions. Cryptography and Communications, 12(5):1011–1033, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00442-
                                        2.

Medina:2020:RHT

[425] Luis A. Medina, Matthew G. Parker, and Pantelimon Stanica. Root-Hadamard transforms and complementary sequences. Cryptography and Communications, 12(5):1035–1049, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00440-
                                        4.

Borissov:2020:NDS

[426] Yuri Borissov and Lyubomir Borissov. A note on the distinctness of some Kloosterman sums. Cryptography and Communications, 12(5):1051–1056, September 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00443-
                                        1.

Ozden:2020:AAS

[427] Büsra Özden and Oguz Yayla. Almost pary sequences. Cryptography and Communications, 12(6):1057–1069, November 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00423-
5.

Jiang:2020:RPT

[428] Jing Jiang and Minquan Cheng. Regular (k, R, 1)-packings with max(R) = 3 and their locally repairable codes. Cryptography and Communications, 12(6):1071– 1089, November 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-020-
00424-4.

Anbar:2020:SPF

[429] Nurdagül Anbar, Canan Kasikçi, and Alev Topuzoglu. Shifted plateaued functions and their differential properties. Cryptography and Communications, 12(6):1091–1105, November 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00426-
2.

Cao:2020:CMP

[430] Yuan Cao, Yonglin Cao, and Fang-Wei Fu. Correcting mistakes in the paper “A mass formula for negacyclic codes of length 2^{k} and some good negacyclic codes over Z_{4} + uZ_{4}” [Cryptogr. Commun. (2017) 9:241– 272]. Cryptography and Communications, 12(6):1107–1110, November 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00429-

z. See [172].

Wang:2020:SCS

[431] Xiaoqiang Wang and Dabin Zheng. The subfield codes of several classes of linear codes. Cryptography and Communications, 12(6):1111–1131, November 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00432-
                                        4.

Beierle:2020:UPN

[432] Christof Beierle and Gregor Leander. 4-uniform permutations with null nonlinearity. Cryptography and Communications, 12(6):1133–1141, November 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00434-
                                        2.

Zheng:2020:SNI

[433] Lijing Zheng, Jie Peng, and Yanjun Li. Several new infinite families of bent functions via second order derivatives. Cryptography and Communications, 12(6):1143–1160, November 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00436-
                                        0.

Calderini:2020:BUS

[434] Marco Calderini and Irene Villa. On the boomerang uniformity of some permutation polynomials. Cryptography and Communications, 12(6):1161– 1178, November 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.

com/article/10.1007/s12095-020-
00439-x.

Gologlu:2020:IFC

[435] Faruk Göloglu. Inverse function is not component-wise uniform. Cryptography and Communications, 12(6):1179– 1194, November 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-020-
00441-3.

Wu:2020:TCA

[436] Xia Wu, Wei Lu, and Xiwang Cao.

Two constructions of asymptotically optimal codebooks via the trace functions. Cryptography and Communications, 12(6):1195–1211, November 2020. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00448-
w.

Fang:2021:EHM

[437] Xiaolei Fang, Meiqing Liu, and Jinquan Luo. On Euclidean hulls of MDS codes. Cryptography and Communications, 13(1):1–14, January 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00428-
0.

Yan:2021:CAC

[438] Ming Yan, Tongjiang Yan, and Yu Li.

Computing the 2-adic complexity of two classes of Ding–Helleseth generalized cyclotomic sequences of periods of twin prime products. Cryptography and Communications, 13(1):15–26, January 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00451-
                                        1.

Sriwirach:2021:RRC

[439] Wateekorn Sriwirach and Chakkrid Klin-eam. Repeated-root constacyclic codes of length 2p^{s} over F_{p^{m}} + uF_{p^{m}} + u^{2}F_{p^{m}}. Cryptography and Communications, 13(1):27–52, January 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00450-
                                        2.

Shao:2021:AMD

[440] Minfeng Shao and Ying Miao. Algebraic manipulation detection codes via highly nonlinear functions. Cryptography and Communications, 13(1):53–69, January 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00453-
                                        z.

Maiti:2021:DFR

[441] Swapan Maiti and Dipanwita Roy Chowdhury. Design of fault-resilient Sboxes for AES-like block ciphers. Cryptography and Communications, 13(1): 71–100, January 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-020-
                                        00452-0.

Ma:2021:SPW

[442] Junru Ma and Jinquan Luo. On symbol-pair weight distribution of MDS codes and simplex codes over finite fields. Cryptography and Communications, 13(1):101–115, January 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00455-
x.

Xiang:2021:TFS

[443] Can Xiang and Wenjuan Yin. Two families of subfield codes with a few weights. Cryptography and Communications, 13(1):117–127, January 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00457-
9.

Aksoy:2021:SDC

[444] Refia Aksoy and Fatma Çaliskan. Selfdual codes over F_{2} × (F_{2} + vF_{2}). Cryptography and Communications, 13(1): 129–141, January 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-020-

00461-z. See correction [457].

Benerjee:2021:CFD

[445] Krishna Gopal Benerjee, Sourav Deb, and Manish K. Gupta. On conflict free DNA codes. Cryptography and Communications, 13(1):143–171, January 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00459-
7.

Wang:2021:CSC

[446] Xiaoqiang Wang, Dabin Zheng, and Yan Zhang. A class of subfield codes of linear codes and their duals. Cryptography and Communications, 13(1):173–196, January 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00460-
0.

Rega:2021:WIT

[447] B. Rega, Z. H. Liu, and C. Durairajan. The t-wise intersection and trellis of relative four-weight codes. Cryptography and Communications, 13(2):197–223, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00456-
w.

Qian:2021:LCO

[448] Liqin Qian, Xiwang Cao, and Sihem Mesnager. Linear codes with one-dimensional hull associated with Gaussian sums. Cryptography and Communications, 13(2):225–243, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00462-
y.

Zhang:2021:FCC

[449] He Zhang and Xiwang Cao. Further constructions of cyclic subspace codes. Cryptography and Communications, 13(2):245–262, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00463-
x.

Bouyukliev:2021:CVW

[450] Iliya Bouyukliev, Stefka Bouyuklieva, and Paskal Piperkov. Characteristic vector and weight distribution of a linear code. Cryptography and Communications, 13(2):263–282, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00458-
8.

Anbar:2021:PPF

[451] Nurdagül Anbar and Canan Kasikci.

Permutations polynomials of the form G(X)^{k} − L(X) and curves over finite fields. Cryptography and Communications, 13(2):283–294, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00465-
9.

Stanica:2021:CDB

[452] Pantelimon Stanica and Aaron Geary.

The c-differential behavior of the inverse function under the EA-equivalence. Cryptography and Communications, 13 (2):295–306, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-020-
00466-8.

Ma:2021:SQC

[453] Fanghui Ma, Jian Gao, and Fang-Wei Fu. (σ, δ)-skew quasi-cyclic codes over the ring Z_{4} + uZ_{4}. Cryptography and Communications, 13(2):307–320, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00467-
7.

Kumar:2021:ARM

[454] Ajeet Kumar, Subhamoy Maitra, and Chandra Sekhar Mukherjee. On approximate real mutually unbiased bases in square dimension. Cryptography and Communications, 13(2):321–329, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-020-00468-
                                        6.

Huczynska:2021:SED

[455] Sophie Huczynska, Christopher Jefferson, and Silvia Nepsinská. Strong external difference families in abelian and non-abelian groups. Cryptography and Communications, 13(2):331–341, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00473-
                                        3.

Zhang:2021:TRB

[456] Jingwei Zhang, Chuangqiang Hu, and Chang-An Zhao. Trace representation of the binary pq^{2}-periodic sequences derived from Euler quotients. Cryptography and Communications, 13(2): 343–359, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00475-1.

Aksoy:2021:CSD

[457] Refia Aksoy and Fatma Çaliskan. Correction to: Self-dual codes over F_{2} × (F_{2} + vF_{2}). Cryptography and Communications, 13(2):361–362, March 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-020-00464-

w. See [444].

Faraj:2021:QPS

[458] Mustafa Faraj and Catherine Gebotys.

Quiescent photonics side channel analysis: Low cost SRAM readout attack. Cryptography and Communications, 13 (3):363–376, May 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-020-
00469-5.

Gologlu:2021:CIS

[459] Faruk Göloglu and Jirí Pavlu.. On CCZ-inequivalence of some families of almost perfect nonlinear functions to permutations. Cryptography and Communications, 13(3):377–391, May 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00476-
0.

Jin:2021:BFS

[460] Wengang Jin, Xiaoni Du, and Cuiling Fan. Boolean functions with six-valued Walsh spectra and their application. Cryptography and Communications, 13 (3):393–405, May 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-021-
00484-0.

He:2021:LCD

[461] Zhiwen He and Jiejing Wen. Linear codes of 2-designs as subcodes of the generalized Reed–Muller codes. Cryptography and Communications, 13(3): 407–423, May 2021. CODEN ????

ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00472-4.

Chathely:2021:CBH

[462] Briji J. Chathely and Rajendra P. Deore. Construction of binary Hadamard codes and their s-PD sets. Cryptography and Communications, 13(3):425–438, May 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00471-
                                        5.

Bartoli:2021:ICM

[463] Daniele Bartoli, Matteo Bonini, and Burçin Günes. An inductive construction of minimal codes. Cryptography and Communications, 13(3):439–449, May 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00474-
                                        2.

Biswas:2021:SRB

[464] Aniruddha Biswas and Palash Sarkar. Separation results for boolean function classes. Cryptography and Communications, 13(3):451–458, May 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00488-
                                        w.

Li:2021:MPI

[465] Yubo Li, Kangquan Li, and Longjiang Qu. More permutations and involutions for constructing bent functions. Cryptography and Communications, 13 (3):459–473, May 2021. CODEN ????

ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-021-
00482-2.

Aubry:2021:SIM

[466] Yves Aubry, Pierre Barthélémy, and Nadia El Mrabet. Special issue from mathematics to embedded devices. Cryptography and Communications, 13(4): 475–477, July 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-021-
00502-1.

Koshelev:2021:HEC

[467] Dmitrii Koshelev. Hashing to elliptic curves of j-invariant 1728. Cryptography and Communications, 13(4):479–494, July 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00478-
y.

Ballet:2021:OSC

[468] Stéphane Ballet, Alexis Bonnecaze, and Thanh-Hung Dang. Optimization of the scalar complexity of Chudnovsky^{2} multiplication algorithms in finite fields. Cryptography and Communications, 13(4):495–517, July 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00494-
y.

Bordage:2021:PCB

[469] Sarah Bordage and Julien Lavauzelle.

On the privacy of a code-based singleserver computational PIR scheme. Cryptography and Communications, 13(4):

519–526, July 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-021-
00477-z.

Cheng:2021:CAL

[470] Wei Cheng, Sylvain Guilley, and Jean-Luc Danger. Categorizing all linear codes of IPM over F_{2^{8}}. Cryptography and Communications, 13(4):527–542, July 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00483-
1.

Landry:2021:MEP

[471] Simon Landry, Yanis Linge, and Emmanuel Prouff. Monomial evaluation of polynomial functions protected by threshold implementations — with an illustration on AES. Cryptography and Communications, 13(4):543–572, July 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00497-
9.

VanOuytsel:2021:HFB

[472] Charles-Henry Bertrand Van Ouytsel, Olivier Bronchain, and François-Xavier Standaert. How to fool a black box machine learning based side-channel security evaluation. Cryptography and Communications, 13(4):573–585, July 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00479-

Momin:2021:SHT

[473] Charles Momin, Olivier Bronchain, and François-Xavier Standaert. A stealthy Hardware Trojan based on a Statistical Fault Attack. Cryptography and Communications, 13(4):587–600, July 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00480-
4.

Dougherty:2021:EGC

[474] S. T. Dougherty, Joe Gildea, and Serap Sahinkaya. G-codes, self-dual G-codes and reversible G-codes over the ring B_{j,k}. Cryptography and Communications, 13(5):601–616, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00487-
x.

Fan:2021:IBS

[475] Yanhong Fan, Sihem Mesnager, and Meiqin Wang. Investigation for 8-bit SKINNY-like S-boxes, analysis and applications. Cryptography and Communications, 13(5):617–636, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00486-
y.

Pasalic:2021:SCM

[476] Enes Pasalic, René Rodríguez, and Yongzhuang Wei. Several classes of minimal binary linear codes violating the Ashikhmin–Barg bound. Cryptography and Communications, 13(5):637– 659, September 2021. CODEN ????

ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00491-1.

Lv:2021:QCC

[477] Jingjie Lv, Ruihu Li, and Yu Yao. Quasicyclic constructions of asymmetric quantum error-correcting codes. Cryptography and Communications, 13(5):661– 680, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00489-9.

Hou:2021:PSF

[478] Xiang dong Hou. A power sum formula by Carlitz and its applications to permutation rational functions of finite fields. Cryptography and Communications, 13(5):681–694, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00495-
                                        x.

Kumar:2021:CLC

[479] Pavan Kumar and Noor Mohammad Khan. A class of linear codes with their complete weight enumerators over finite fields. Cryptography and Communications, 13(5):695–725, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00496-
                                        w.

Shi:2021:QSC

[480] Xiaoping Shi, Qin Yue, and Xinmei Huang. Quantum synchronizable codes from the Whiteman’s generalized cyclotomy. Cryptography and Communications, 13(5):727–739, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00501-
2.

Meaux:2021:FAI

[481] Pierrick Méaux. On the fast algebraic immunity of threshold functions. Cryptography and Communications, 13(5):741–762, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00505-
y.

Liu:2021:TPC

[482] Chen Liu, Shuaijun Liu, and Zhengchun Zhou. Three-phase Z-complementary triads and almost complementary triads. Cryptography and Communications, 13(5):763–773, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00509-
8.

Arasu:2021:NNR

[483] K. T. Arasu, Daniel M. Gordon, and Yiran Zhang. New nonexistence results on circulant weighing matrices. Cryptography and Communications, 13(5):775–789, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00492-

Jamet:2021:MOC

[484] Damien Jamet, Pierre Popoli, and Thomas Stoll. Maximum order complexity of the sum of digits function in Zeckendorf base and polynomial subsequences. Cryptography and Communications, 13(5):791–814, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00507-
                                        w.

Kolsch:2021:FSD

[485] Lukas Kölsch and Robert Schüler. Formal self duality. Cryptography and Communications, 13(5):815–836, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00508-
                                        9.

Otal:2021:EGQ

[486] Kamil Otal and Eda Tekin. Efficient generation of quadratic cyclotomic classes for shortest quadratic decompositions of polynomials. Cryptography and Communications, 13(5):837– 845, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00512-z.

Yao:2021:CCM

[487] Ge Yao and Udaya Parampalli. Cryptanalysis of the class of maximum period Galois NLFSR-based stream ciphers. Cryptography and Communications, 13(5):847–864, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00511-
0.

Xiao:2021:ACT

[488] Zibi Xiao and Xiangyong Zeng. 2adic complexity of two constructions of binary sequences with period 4N and optimal autocorrelation magnitude. Cryptography and Communications, 13(5):865–885, September 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00498-
8.

Budaghyan:2021:ESI

[489] Lilya Budaghyan, Claude Carlet, and Kaisa Nyberg. Editorial: Special issue on Boolean functions and their applications 2020. Cryptography and Communications, 13(6):887–889, November 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00536-
5.

Stanica:2021:CWS

[490] Pantelimon Stanica, Constanza Riera, and Anton Tkachenko. Characters, Weil sums and c-differential uniformity with an application to the perturbed Gold function. Cryptography and Communications, 13(6):891–907, November 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00485-
z.

Kolomeec:2021:SGP

[491] Nikolay Kolomeec. Some general properties of modified bent functions through addition of indicator functions. Cryptography and Communications, 13(6):909–926, November 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00528-
                                        5.

Ding:2021:LCE

[492] Cunsheng Ding and Chunming Tang. The linear codes of t-designs held in the Reed–Muller and Simplex codes. Cryptography and Communications, 13(6):927–949, November 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00470-
                                        6.

Mesnager:2021:CWP

[493] Sihem Mesnager and Sihong Su. On constructions of weightwise perfectly balanced Boolean functions. Cryptography and Communications, 13(6):951– 979, November 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00481-3.

Chase:2021:KTA

[494] Benjamin Chase and Petr Lisonek. Kimtype APN functions are affine equivalent to Gold functions. Cryptography and Communications, 13(6):981– 993, November 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.

com/article/10.1007/s12095-021-
00490-2.

Kaleyski:2021:IEC

[495] Nikolay S. Kaleyski. Invariants for EA- and CCZ-equivalence of APN and AB functions. Cryptography and Communications, 13(6):995–1023, November 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00541-
8.

Ding:2021:SSS

[496] Jian Ding, Changlu Lin, and Sihem Mesnager. Secret sharing schemes based on the dual of Golay codes. Cryptography and Communications, 13(6):1025– 1041, November 2021. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-021-
00531-w.

Wang:2022:FOT

[497] Dandan Wang and Xiwang Cao.

A family of optimal ternary cyclic codes with minimum distance five and their duals. Cryptography and Communications, 14(1):1–13, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00493-
z.

Allahmadi:2022:NCE

[498] A. Allahmadi, A. AlKenani, and P. Solé.

New constructions of entanglementassisted quantum codes. Cryptography and Communications, 14(1):15–37, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00499-

7. See correction [499].

Allahmadi:2022:CNC

[499] A. Allahmadi, A. AlKenani, and P. Solé. Correction to: New constructions of entanglement-assisted quantum codes. Cryptography and Communications, 14 (1):39, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-

00532-9. See [498].

Dhooghe:2022:RUA

[500] Siemen Dhooghe and Svetla Nikova. Resilient uniformity: applying resiliency in masking. Cryptography and Communications, 14(1):41–58, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00515-
                                        w.

Islam:2022:CLN

[501] Habibul Islam and Om Prakash. Construction of LCD and new quantum codes from cyclic codes over a finite non-chain ring. Cryptography and Communications, 14(1):59–73, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00516-
                                        9.

Kim:2022:PEP

[502] Kwang Ho Kim, Sihem Mesnager, and Dok Nam Lee. Preimages of p-linearized polynomials over F_{p}. Cryptography and Communications, 14(1):75–86, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00514-
x.

Lao:2022:NCD

[503] Huimin Lao, Hao Chen, and Xiaoqing Tan. New constant dimension subspace codes from block inserting constructions. Cryptography and Communications, 14(1):87–99, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00524-
9.

Kudin:2022:PLS

[504] S. Kudin, E. Pasalic, and F. Zhang.

Permutations without linear structures inducing bent functions outside the completed Maiorana–McFarland class. Cryptography and Communications, 14 (1):101–116, January 2022. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL https:

//link.springer.com/article/10.
1007/s12095-021-00523-w.

Anupindi:2022:LCS

[505] Vishnupriya Anupindi and László Mérai.

Linear complexity of some sequences derived from hyperelliptic curves of genus 2. Cryptography and Communications, 14(1):117–134, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00521-
y.

Gu:2022:NFP

[506] Zhi Gu, Zhengchun Zhou, and Udaya Parampalli. A new family of polyphase sequences with low correlation. Cryptography and Communications, 14(1): 135–144, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00522-x.

Fang:2022:NGH

[507] Xiaolei Fang, Renjie Jin, and Wen Ma. New Galois hulls of GRS codes and application to EAQECCs. Cryptography and Communications, 14(1):145–159, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00525-
                                        8.

Wan:2022:EBT

[508] Qianhong Wan and Chao Li. On equivalence between two known families of APN polynomial functions and APN power functions. Cryptography and Communications, 14(1):161–182, January 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00533-
                                        8.

Edemskiy:2022:SAC

[509] Vladimir Edemskiy and Yuhua Sun. The symmetric 2-adic complexity of sequences with optimal autocorrelation magnitude and length 8q. Cryptography and Communications, 14(2):183–199, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s12095-021-00503-
0.

Niu:2022:NCD

[510] Yongfeng Niu, Qin Yue, and Daitao Huang. New constant dimension subspace codes from parallel linkage construction and multilevel construction. Cryptography and Communications, 14(2):201–214, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00504-
z.

Liu:2022:CAI

[511] Yan Liu and Xiwang Cao. A class of affine-invariant codes and their support 2-designs. Cryptography and Communications, 14(2):215–227, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00506-
x.

Yao:2022:ITA

[512] Ge Yao and Udaya Parampalli. Improved transformation algorithms for generalized Galois NLFSRs. Cryptography and Communications, 14(2): 229–258, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-021-
00500-3.

Xu:2022:HLC

[513] Heqian Xu and Wei Du. Hermitian LCD codes over F_{q^{2}} + uF_{q^{2}} and their applications to maximal entanglement EAQECCs. Cryptography and Communications, 14(2):259–269, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00510-
                                        1.

Kaleyski:2022:DEE

[514] Nikolay Kaleyski. Deciding EAequivalence via invariants. Cryptography and Communications, 14(2):271–290, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00513-
                                        y.

Zhou:2022:COL

[515] Limengnan Zhou, Xing Liu, and Changyuan Wang. Classes of optimal low-hit-zone frequency-hopping sequence sets with new parameters. Cryptography and Communications, 14(2): 291–306, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00518-7.

Zhou:2022:MPW

[516] Limengnan Zhou and Hanzhou Wu. Multi-party watermark embedding with frequency-hopping sequences. Cryptography and Communications, 14(2): 307–318, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00517-8.

Budaghyan:2022:BSA

[517] Lilya Budaghyan, Nikolay Kaleyski, and Pantelimon Stanica. On the behavior of some APN permutations under swapping points. Cryptography and Communications, 14(2):319–345, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00520-
z.

Liu:2022:NFH

[518] Xing Liu, Shaofang Hong, and Limengnan Zhou. NHZ frequency hopping sequence sets under aperiodic Hamming correlation: Tighter bound and optimal constructions. Cryptography and Communications, 14(2):347–356, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00527-
6.

Yan:2022:DUT

[519] Haode Yan. On (−1)-differential uniformity of ternary APN power functions. Cryptography and Communications, 14(2):357–369, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00526-
7.

Jiang:2022:DBS

[520] Sha Jiang, Kangquan Li, and Longjiang Qu. Differential and boomerang spectrums of some power permutations. Cryptography and Communications, 14 (2):371–393, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-021-
00530-x.

Edemskiy:2022:ELC

[521] Vladimir Edemskiy and Nikita Sokolovskiy. The estimate of the linear complexity of generalized cyclotomic binary and quaternary sequences with periods p^{n} and 2p^{n}. Cryptography and Communications, 14(2):395–414, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00534-
                                        7.

Xu:2022:MLC

[522] Guangkui Xu, Longjiang Qu, and Gaojun Luo. Minimal linear codes from weakly regular bent functions. Cryptography and Communications, 14(2): 415–431, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00519-6.

Vielhaber:2022:RCB

[523] Michael Vielhaber, Mónica del Pilar Canales Chacón, and Sergio Jara Ceballos. Rational complexity of binary sequences, FQSRs, and pseudoultrametric continued fractions in R. Cryptography and Communications, 14 (2):433–457, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00539-2.

Salagean:2022:IBP

[524] Ana Salagean and Pantelimon Stanica. Improving bounds on probabilistic affine tests to estimate the nonlinearity of Boolean functions. Cryptography and Communications, 14(2):459–481, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00529-
4.

Li:2022:JHR

[525] Haitao Li, Yang Liu, and Gang Wang.

Jump and hop randomness tests for binary sequences. Cryptography and Communications, 14(2):483–502, March 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00538-
3.

Barbero:2022:PSI

[526] Ángela Barbero, Vitaly Skachek, and Øyvind Ytrehus. Preface of special issue on coding theory and applications. Cryptography and Communications, 14 (3):503, May 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-022-
00560-z.

Gutierrez:2022:ALC

[527] Jaime Gutierrez. Attacking the linear congruential generator on elliptic curves via lattice techniques. Cryptography and Communications, 14(3):505–525, May 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00535-
6.

Kumar:2022:RBD

[528] Ajeet Kumar and Subhamoy Maitra.

Resolvable block designs in construction of approximate real MUBs that are sparse. Cryptography and Communications, 14(3):527–549, May 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00537-
                                        4.

Kumar:2022:RRC

[529] Raj Kumar and Maheshanand Bhaintwal. Repeated root cyclic codes over Z_{p^{2}} + uZ_{p^{2}} and their Lee distances. Cryptography and Communications, 14 (3):551–577, May 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00540-9.

Li:2022:HDC

[530] Fengwei Li. The Hermitian dualcontaining LCD BCH codes and related quantum codes. Cryptography and Communications, 14(3):579–596, May 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-021-00543-
                                        6.

Liang:2022:CAH

[531] Huicong Liang, Sihem Mesnager, and Meiqin Wang. Cryptanalysis of the AEAD and hash algorithm DryGAS-CON. Cryptography and Communications, 14(3):597–625, May 2022. CO-DEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s12095-021-00542-7.

Shi:2022:LAC

[532] Minjia Shi, Shitao Li, and Patrick Solé. LCD and ACD codes over a noncommutative non-unital ring with four elements. Cryptography and Communications, 14(3):627–640, May 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00545-
4.

Xiang:2022:SDB

[533] Can Xiang. Some t-designs from BCH codes. Cryptography and Communications, 14(3):641–652, May 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00546-
3.

Li:2022:SOZ

[534] Xia Li, Qin Yue, and Deng Tang.

The second-order zero differential spectra of almost perfect nonlinear functions and the inverse function in odd characteristic. Cryptography and Communications, 14(3):653–662, May 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00544-
5.

Liu:2022:CBC

[535] Kaiqiang Liu, Qi Wang, and Haode Yan.

A class of binary cyclic codes with optimal parameters. Cryptography and Communications, 14(3):663–675, May 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00548-
1.

Fang:2022:NCS

[536] Weijun Fang, Jun Zhang, and Fang-Wei Fu. New constructions of selfdual generalized Reed–Solomon codes. Cryptography and Communications, 14 (3):677–690, May 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00549-0.

Mesnager:2022:SRT

[537] Sihem Mesnager, Bimal Mandal, and Mounira Msahli. Survey on recent trends towards generalized differential and boomerang uniformities. Cryptography and Communications, 14(4): 691–735, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-021-
                                        00551-6.

Meidl:2022:SAG

[538] Wilfried Meidl. A survey on p-ary and generalized bent functions. Cryptography and Communications, 14(4): 737–782, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-022-
                                        00570-x.

Merai:2022:PSD

[539] László Mérai and Arne Winterhof. Pseudorandom sequences derived from automatic sequences. Cryptography and Communications, 14(4):783–815, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-022-00556-

Shen:2022:NCC

[540] Bingsheng Shen, Yang Yang, and Zhengchun Zhou. New z-complementary/ complementary sequence sets with nonpower-of-two length and low PAPR. Cryptography and Communications, 14 (4):817–832, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s12095-021-
00550-7.

Araya:2022:CQO

[541] Makoto Araya and Masaaki Harada. On the classification of quaternary optimal Hermitian LCD codes. Cryptography and Communications, 14(4):833–847, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00552-
5.

Jeon:2022:DUL

[542] Yongjin Jeon, Seungjun Baek, and Jongsung Kim. Differential uniformity and linearity of S-boxes by multiplicative complexity. Cryptography and Communications, 14(4):849–874, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-021-00547-
2.

Wu:2022:MLC

[543] Xia Wu, Wei Lu, and Xiwang Cao.

Minimal linear codes constructed from functions. Cryptography and Communications, 14(4):875–895, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s12095-021-00553-
                                        4.

Figueiredo:2022:TCM

[544] Ravi B. D. Figueiredo and Juliano B. Lima. Tangent-Chebyshev maps over finite fields: New properties and functional graphs. Cryptography and Communications, 14(4):897–908, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-022-00565-
                                        8.

Dinh:2022:SQC

[545] Hai Q. Dinh, Narendra Kumar, and Abhay Kumar Singh. A study of quantum codes obtained from cyclic codes over a non-chain ring. Cryptography and Communications, 14(4):909–923, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s12095-022-00567-
                                        6.

Bartoli:2022:CAP

[546] Daniele Bartoli and Marco Timpanella. On a conjecture on APN permutations. Cryptography and Communications, 14 (4):925–931, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s12095-022-
                                        00558-7.

Kose:2022:FSP

[547] Seyda Köse and Ferruh Özbudak. Factorization of some polynomials over finite local commutative rings and applications to certain self-dual and LCD codes. Cryptography and Communications, 14(4):933–948, July 2022. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic).

URL https://link.springer.com/
article/10.1007/s12095-022-00557-